### v1

# EDEN: Efficient Dual-Layer Exploration Planning for Fast UAV Autonomous Exploration in Large 3-D Environments

Abstract
Efficient autonomous exploration in large-scale environments remains challenging due to the high planning computational cost and low-speed maneuvers. In this paper, we propose a fast and computation-efficient dual-layer exploration planning method. The insight of our dual-layer method is efficiently finding an acceptable long-term region routing and greedily exploring the target in the region of the first routing area with high speed. Specifically, the proposed method finds the long-term area routing through an approximate algorithm to ensure real-time planning in large-scale environments. Then, the viewpoint in the first routing region with the lowest curvature-penalized cost which can effectively reduce decelerations caused by sharp turn motions will be chosen as the next exploration target. To further speed up the exploration, we adopt an aggressive and safe exploration-oriented trajectory to enhance exploration continuity. The proposed method is compared to state-of-the-art methods in challenging simulation environments. The results show the proposed method outperforms other methods in terms of exploration efficiency, computational cost, and trajectory speed. We also conduct real-world experiments to validate the effectiveness of the proposed method. The code will be open-sourced [^1] https://github.com/NKU-MobFly-Robotics/EDEN.
Index Terms:
Autonomous exploration, motion planning, trajectory planning.
I Introduction
Umanned aerial vehicles (UAVs) are popular in many 3-D exploration relevant applications, e.g., search and rescue [1, 2], environment inspection [3, 4], environment reconstruction [5, 6, 7, 8], among others. However, most existing methods remain low efficiency, low trajectory speed, and high computational cost.
Existing exploration methods can be coarsely separated into two kinds of methods: greedy methods and optimal-routing-based methods. Greedy methods attempt to select the target that maximizes the reward in a short period. The original frontier-based method [9] selects the closest frontier, i.e., the boundary voxel of explored and unknown space as an exploration target. To improve the UAV speed, [10] propose to select a frontier in camera field-of-view (FOV) to minimize the velocity change. By combining Next-Best-Views and rapid random tree (RRT), receding-horizon-next-best-views (RH-NBVP) [11] is proposed, and its variant methods [12, 13, 14, 15] show high exploration efficiency in small environments. However, these greedy methods suffer from massive back-and-forth maneuvers in large-scale environments due to the lack of long-term consideration. Another type of method, optimal-routing-based methods [16, 17, 18, 19, 20, 21], make decisions by constructing and solving routing problems to minimize the total routing length. The work in [21] proposes to find a coverage path by solving the traveling salesman problem (TSP). The method proposed by [16] constructs and solves the asymmetrical traveling salesman problem (ATSP) to find the shortest routing of frontiers. To further reduce back-and-forth maneuvers, [18] increases the routing priority of isolated frontiers. By solving routing problems, these methods dramatically reduce back-and-forth maneuvers. However, the computational cost grows fast with the number of routing nodes, which makes these methods have to execute stops for planning in large-scale environments. What is more, the work in [22] shows that the speed of UAVs in exploration still has great potential to be stimulated. Therefore, a computation-efficient, far-sighted, and high-speed method is expected.
Motivated by the above drawbacks of existing methods, we propose an Efficient Dual-layer Exploration plaNning method (EDEN) which can plan efficient exploration motions in hundred-meter-level 3-D environments in real-time. Different from greedy methods and optimal-routing-based methods, the proposed dual-layer method first finds a long-term sub-optimal exploration region routing efficiently and then decides the short-term targets that minimizes the exploration motion cost greedily. Specifically, we improve the Christofides algorithm, which is an approximate algorithm for solving TSP, to an exploration-oriented heuristic Christofides (EOHC) algorithm to find a sub-optimal but satisfying tour for ATSP with low computational cost. Then, a viewpoint for safety and two viewpoints with the highest curvature-penalized exploration gain in the first exploration region of the routing will be selected for high-speed trajectory planning. Then, inspired by [23], we propose an aggressive and safe exploration-oriented (ASEO) trajectory for high-speed continuous exploration. The ASEO trajectory is composed of three parts: an exploring part, a continuous part, and a safety part. The time of the exploring part, the time of the continuous part, and the feasibility of the safety part are optimized together to find an aggresive trajectory and a backup trajectory for safety.
We evaluate the proposed method by comparing it against state-of-the-art methods in various challenging environments. The results show that the proposed method has $5.2\%-14.0\%$ higher exploration efficiency, $10.4\%-15.5\%$ faster UAV speed, and $28.9\%-74.5\%$ lower computational cost. Further, the proposed method is deployed on a UAV and is tested in real-world environments to evaluate its effectiveness. The contributions of this work are summarized as follows:
  1. 1.
    We propose an exploration-oriented heuristic Christofides (EOHC) algorithm for efficiently solving large-scale ATSP. The proposed EOHC can find a satisfactory region routing with a low computational cost. The real-time performance of EOHC can be guaranteed even in hundred-meter-level environments.
  2. 2.
    To take advantage of the current UAV speed and enhance exploration continuity, we propose a heuristic curvature-penalized continuous gain for the short-term exploration targets choosing.
  3. 3.
    To make the UAV keep high-speed continuous exploration in unknown environments, we propose an ASEO trajectory planning method that ensures trajectory efficiency and safety at the same time.
  4. 4.
    Massive comparative simulations are conducted to evaluate the proposed method. We also test the proposed method by real-world experiments. We will release our source code.
II related work
II-A Greedy Exploration Planning
The frontier-based method is first proposed by Yamauchi et al. [9]. It selects the frontier voxel that is closest to the robot as the exploration target. To minimize the UAV velocity change, [10] selects the frontier voxel that is located at the boundary of camera FOV as the exploration target. Different from the greedy frontier-based methods that only consider the next step exploration, Bircher et al. [11] propose RH-NBVP which samples a set of exploration motion sequences and executes the most informative sequence’s first step and it shows high efficiency in short-term exploration. Nevertheless, in large-scale environments, it suffers from sample failure. To cope with this problem, Selin et al. [24] use frontier to guide sampling when a local sample fails. However, long-distance sampling is time-consuming and it does not guarantee success even with the guidance of frontier. To reduce the planning cost and improve long-distance exploration, roadgraphs, which can compress the complex explorable space information and tarversable space information into compact nodes and edges, are widely used in recent proposed greedy methods [12, 25, 14, 26, 27, 28, 29, 30]. Consequently, roadgraphs enable greedy methods of efficient planning in large environments. Although these methods show high efficiency in short-term exploration and low computational cost, they suffer from back-and-forth maneuvers due to a lack of long-term consideration.
II-B Optimal-Routing-Based Exploration Planning
Optimal-routing-based methods usually formulate the exploration target decision problems into routing problems. Optimal-tour-based methods can be roughly devided into frontier-ATSP methods [16, 17, 31, 18] and coverage-path-ATSP methods [32, 20, 21, 5, 19]. Frontier-ATSP methods formulate ATSP using frontiers and find the shortest tour between frontiers. Coverage-path-ATSP methods formulate ATSP with unexplored space and find the shortest coverage path. The main drawback of the optimal-routing-based methods is that they have to spend luxury computation on solving ATSP in large-scale environments. Moreover, counterintuitively, recent works [28, 33] report that optimal-tour-based methods usually do not have an obvious advantage in minimizing the exploration path distance compared to greedy methods. This is because optimal-routing-based methods have implicit optimal assumptions, where frontier-ATSP methods assume a new frontier will not be generated after exploring an existing frontier, while coverage-path-ATSP methods assume all the unexplored space is traversable. Obviously, the above assumptions do not hold in most cases. Therefore, spending too much computation on finding a high-quality ATSP solution is not cost-effective.
II-C Exploration-Oriented Trajectory Planning
Exploration trajectory plays a crucial role in exploration tasks. The early proposed exploration methods [11, 9] stop-and-go motions are used. To utilize the good maneuverabilities of UAVs, Dharmadhikari et, al. [15] adopts motion primitives to speed up the exploration. To further improve the trajectory speed, Zhou et al. [16] optimizes minimum-time trajectory with optimizable terminal velocity and acceleration for high-speed exploration. Although minimum-time trajectory is very efficient, the optimizable terminal velocity and acceleration do not ensure being beneficial to continuously exploring the newly generated frontier after exploring the previous frontier. What is more, the optimizable terminal velocity and acceleration can lead to collision.
In summary, a far-sighted and computationally efficient exploration planning method and a trajectory planning method that enable fast continuous exploration while ensuring safety are required for exploration in large-scale environments.
Figure 1: The overview of the proposed method framework.
III system overview
The proposed method is composed of a dual-layer exploration target decision module that can plan both the long-term region routing and the short-term viewpoint target, and a trajectory planning module that can plan high-speed continuous exploration trajectories as shown in Fig. 1. A dynamic voxel map with resolution $r_{v}$ is updated by integrating depth image and odometry information. Different from the normal voxel map, the dynamic voxel map only maintains a local part of voxels around the UAV while other voxels are dynamically read and written to save memory usage in large environments. Based on the voxel map, a dynamic topological graph (DTG), is maintained to manage exploration-related information. The traversable information is maintained by DTG’s history nodes and edges between them, and the explorable space information is maintained by DTG’s explorable region of interest (EROI), which is a set of uniform scale cubic regions that can be explored by visiting corresponding active viewpoints. For detailed information, we refer readers to our previous work [34]. The exploration planning utilizes DTG to efficiently formulate ATSP, and ATSP is solved through EOHC to find satisfactory exploration region routing with small computation. Then, three viewpoints, including an exploring viewpoint, a continuous viewpoint, and a safety viewpoint in the first region of the region routing are found for the following trajectory planning. Then, an aggressive and safe exploration-oriented trajectory which separates at the exploring viewpoint, and its two branches end at a continuous viewpoint and a safe viewpoint correspondingly, is optimized. The exploring part of the aggressive and safe exploration-oriented trajectory will be executed.
IV Dual-layer exploration viewpoints planning
In this section, we first propose the EOHC algorithm for finding a satisfactory region routings efficiently in large-scale environments. Then, we use curvature-penalized continuous gain to decide the targets in the first region of the tour for the following high-speed ASEO trajectory planning.
Figure 2: An instance of EOHC-based routing planning, the numbers indicate the lengths of edges. (a) An MST that connects all the explorable history nodes of DTG is extracted. (b) The leaf point, whose distance is the longest to the root node, and its corresponding branch (green) will be executed last. (c) Once, there are multiple siblings for the current node, the sibling (orange dashed curve) whose first node has the shortest distance to the current node will be chosen as the next branch. (d) The final routing found by the proposed EOHC-based routing planning.
IV-A EOHC-Based Global Routing Planning
As analyzed in Section. II.B, the optimal routing of frontiers or unexplored space is very likely not the optimal routing for exploration due to the implicit optimal assumption not being met. Therefore, instead of finding a high-quality of all precise explorable frontiers, the proposed method aims to spend low computation on finding a satisfactory coarse routing of explorable regions. The process of EOHC-based routing planning is shown in Fig. 2.
Firstly, we extract all the history nodes of DTG that connect at least one EROI as the explorable regions. To reduce the computation when formulating ATSP, the distances between all explorable regions and the distances between the robot and all explorable regions are computed by conducting the Dijkstra algorithm on DTG. Thanks to the sparse DTG, the ATSP can be formulated efficiently.
To solve ATSP efficiently, we design an EOHC algorithm, which is improved from the Christofides algorithm [35] and tailored to ATSP. The key idea of the Christofides algorithm is to construct a minimum spanning tree (MST, Fig. 2(a)) using the Prim algorithm and find a walk that goes around the MST. The solution routing of the Christofides algorithm ensures no longer than twice the optimal routing. However, the Christofides algorithm is designed for symmetric TSP. For exploration-oriented ATSP, the robot does not have to return to the original node as stated by [16]. Therefore, to minimize the routing length, the leaf node that is the farthest from the robot and nodes on its corresponding branch (in Fig. 2(b), the green nodes) will be executed last to save the highest cost of returning to the original node. Another problem is how to decide the sequence of other branches. Since all the regions will be explored finally, the proposed method chooses the next node closest to the current node heuristically when the current node has multiple siblings as shown in Fig. 2(c). Finally, a routing for ATSP will be found as shown in Fig. 2(d).
For the worst case, the path length $l_{eohc}$ of the routing solved by the EOHC algorithm holds:
$l_{eohc}\leq 2l^{*}+l_{m},$ (1)
where $l^{*}$ is the optimal solution of the corresponding ATSP, and $l_{m}$ is the longest distance among all the distances between the robot and all the explorable history nodes. The computational complexity of the proposed EOHC is $\mathcal{O}(n^{3})$, where $n$ is the total number of nodes.
IV-B Curvature-Penalized Viewpoints Planning
After finding the long-term routing of regions, three viewpoints (an exploring viewpoint, a continuous viewpoint, and a safety viewpoint) in the first region of the routing will be selected for ASEO trajectory planning. The UAV will go to explore the exploring viewpoint (in the explored map) and maintain a speed and an acceleration conducive to explore the continuous viewpoint (in the explored map or unexplored map) if no obstacle is detected, otherwise it can return to the safety viewpoint (in the explored map).
For detailed viewpoint evaluation, existing methods usually use the path distance and turning angle to avoid long traveling distances and sharp turns in a decoupled manner. However, the maximum velocity is limited by the curvature of the trajectory at a point according to the circular motion model. We compute the velocity $v_{s}(\mathbf{p})$ at $\mathbf{p}$ by
$v_{s}(\mathbf{p})=\min\left(\sqrt{\frac{\lambda_{a}\cdot a_{m}}{K_{c}(\textbf{p})}},v_{m}\right),$ (2)
where $v_{m}$ is the maximum linear velocity of the UAV, $a_{m}$ is the maximum linear acceleration of the UAV, $K_{c}(\mathbf{p})$ is the curvature of the trajectory at $\mathbf{p}$, and $\lambda_{a}\in(0.0,1.0)$ is an adjustable parameter. Based on this fact, we designed a curvature-penalized heuristic motion cost evaluation method to select the best exploring target and the best continuous target for high-speed exploration.
IV-B1 Curvature-Penalized Exploring Viewpoint and Continuous Viewpoint Planning
An exploring viewpoint in the first region of routing and a continuous viewpoint located in the corresponding EROI of the exploring viewpoint will be selected for aggressive high-speed trajectory planning.
First, a Dijkstra search starts from the current UAV position in the explored space is conducted to find the path to each candidate viewpoint in the first region of the routing. Then, for each path to a candidate viewpoint, an anchor point on the path which is $\lambda_{arc}\cdot v_{m}$ ($\lambda_{arc}$ is a tunable parameter) away from the UAV position is selected to formulate a virtual circular arc trajectory. The formulation of the $i$-th virtual circular arc trajectory to the $k$-th candidate viewpoint $\textbf{vp}_{k}=\left\{\mathbf{p}_{k},\psi_{k}\right\}\in\mathbb{R}^{3}\times SO(2)$ is constrained by the current UAV position $\mathbf{p}_{c}\in\mathbb{R}^{3}$, the current UAV speed $\dot{\mathbf{p}}_{c}\in\mathbb{R}^{3}$, and the anchor point $\mathbf{p}_{a,k}\in\mathbb{R}^{3}$. The center point of the circular arc $\mathbf{p}_{ac,k}\in\mathbb{R}^{3}$ and its curvature can be computed by
$\mathbf{p}_{ac,k}=$
[…]
$K_{c,k}=\frac{1}{\|\mathbf{p}_{ac,k}-\mathbf{p}_{c}\|_{2}}.$ (3)
The virtual velocity $v_{as,k}$ to $\textbf{vp}_{k}$ is penalized by the curvature $K_{c,k}$ and is computed by Eq. 2. The motion cost of going to this exploring target is computed by
$\mathrm{CostE}(\mathbf{vp}_{k})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{c})}{v_{as,k}},\frac{\mathrm{YD}(\psi_{k},\psi_{c})}{\dot{\psi}_{m}}\right),$ (4)
where the function $\mathrm{PL}(\mathbf{p}_{c}$, $\mathbf{p}_{k})$ computes the length of the path from $\mathbf{p}_{c}$ to $\mathbf{p}_{k}$, $\mathrm{YD}(\psi_{k},\psi_{c})$ the yaw difference between the current UAV yaw $\psi_{c}$ and $\psi_{k}$, and $\dot{\psi}_{m}$ is the maximum yaw velocity. Since the terminal velocity at $\mathbf{vp}_{k}$ is unknown, we continue to use the $v_{as,k}$ for computing the continuous motion cost from $\mathbf{vp}_{k}$ to the $j$-th viewpoint $\mathbf{vp}_{j}=\left\{\mathbf{p}_{j},\psi_{j}\right\}\in\mathbb{R}^{3}\times SO(2)$ inside the EROI of $\mathbf{vp}_{k}$
$\mathrm{CostC}(\mathbf{vp}_{j})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{j})}{v_{as,k}},\frac{\mathrm{YD}(\psi_{j},\psi_{k})}{\dot{\psi}_{m}}\right).$ (5)
Finnally, the gain of this viewpoint-pair is evaluated by
$\mathrm{G}(\mathbf{vp}_{k},\mathbf{vp}_{j})=\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{k})}\cdot(1+\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{j})}).$ (6)
Here, $\lambda_{e}$ is a tunable parameter. The larger $\lambda_{e}$ is, the more sensitive it is to the motion costs. The viewpoint-pair with the highest gain is selected as the exploring viewpoint $\mathbf{vp}_{e}$ and the continuous viewpoint $\mathbf{vp}_{c}$.
It is worth noting that the voxels on the continuous line between $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ can not be occupied but free and unknown. This ensures the existence of the possibility for continuous exploration. To efficiently check the feasibilities of the viewpoint pairs, we propose a continuous line-checking method that follows the fast collision checking fashion [36] to avoid recalculation of indexes of voxels at overlapping positions of continuous lines. Specifically, our method generates a table of indexes of voxel that needs to be checked in advance and iterates the table when checking an EROI. Once an iterative voxel is found occupied, its corresponding viewpoint pairs are confirmed infeasible and will not be considered.
IV-B2 Safety Viewpoint Planning
After finding $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$, another viewpoint in the first region of routing will be selected. The safety viewpoint $\mathbf{vp}_{s}$ is the backup target if new obstacles are detected on the trajectory between $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$. To find $\mathbf{vp}_{c}$ efficiently, we directly conduct the Dijkstra algorithm starting from $\mathbf{vp}_{e}$, and the viewpoint that closest to $\mathbf{vp}_{e}$ will be selected as $\mathbf{vp}_{c}$.
Figure 3: An instance of ASEO trajectory. The initial position, velocity, and acceleration of $\mathcal{J}_{e}$ (violet) are fixed with the current UAV position, velocity, and acceleration (yellow). The continuous trajectory $\mathcal{J}_{c}$ (red) ends at $\mathbf{vp}_{c}$. The safety trajectory $\mathcal{J}_{s}$ (red) ends at $\mathbf{vp}_{s}$. The optimizable parameters, including intermediate 4-D points, trajectory durations, velocity, and acceleration at $\mathbf{vp}_{e}$, are illustrated in green.
IV-C ASEO Trajectory Optimization
Inspired by [23], we propose to optimize our 4-D aggressive and safe exploration-oriented (ASEO) trajectory which is based on MINCO trajectory [37]. Different from the 3-D MINCO trajectory, the fourth dimension is the yaw angle, and we optimize it with the position trajectory together.
A 4-D ASEO trajectory $\mathcal{J}(t):\left[t_{s},t_{e}\right]\mapsto\mathbb{R}^{3}\times SO(2)$ ($t_{s}$ and $t_{e}$ are the start time and end time of the trajectory) is paramerized by the initial state $\textbf{S}_{i}=\left\{\mathbf{vp}_{i},\dot{\mathbf{vp}}_{i},\ddot{\mathbf{vp}}_{i}\right\}$, the terminal state $\mathbf{S}_{t}=\left\{\mathbf{vp}_{t},\dot{\mathbf{vp}}_{t},\ddot{\mathbf{vp}}_{t}\right\}$, the $m-1$ intermediate viewpoints $\mathbf{VP}=\left\{\mathbf{vp}_{1},\mathbf{vp}_{2},\cdots,\mathbf{vp}_{m-1}\right\}$, and $m$ durations of trajectory segments $\Delta\mathbf{T}=\left\{\Delta t_{1},\Delta t_{2},\cdots,\Delta t_{m}\right\}$. As shown by Fig. 3, the proposed ASEO trajectory is composed of three connected MINCO trajectories: an exploring trajectory $\mathcal{J}_{e}$ (paramerized by $\textbf{S}_{i,e}$, $\textbf{S}_{t,e}$, $\mathbf{VP}_{e}$, and $\Delta\mathbf{T}_{e}$), a continuous trajectory $\mathcal{J}_{c}$ (paramerized by $\textbf{S}_{i,c}$, $\textbf{S}_{t,c}$, $\mathbf{VP}_{c}$, and $\Delta\mathbf{T}_{c}$), and a safety trajectory $\mathcal{J}_{s}$ (paramerized by $\textbf{S}_{i,s}$, $\textbf{S}_{t,s}$, $\mathbf{VP}_{s}$, and $\Delta\mathbf{T}_{s}$). Some parameters of the ASEO trajectory are fixed. The exploring trajectory starts from the current UAV position $\{\mathbf{p}_{c},\psi_{c}\}$, velocity $\{\dot{\mathbf{p}}_{c},\dot{\psi}_{c}\}$, and acceleration $\{\ddot{\mathbf{p}}_{c},\ddot{\psi}_{c}\}$. The three trajectories are connected at the exploring viewpoint $\mathbf{vp}_{e}$. The continuous trajectory and the safety trajectory end at $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$ correspondingly, their velocity and acceleration are set by zero. We summarize these fixed parameters as:
$\mathbf{vp}_{i,e},\dot{\mathbf{vp}}_{i,e},\ddot{\mathbf{vp}}_{i,e}=\{\mathbf{p}_{c},\psi_{c}\},\{\dot{\mathbf{p}}_{c},\dot{\psi}_{c}\},\{\ddot{\mathbf{p}}_{c},\ddot{\psi}_{c}\},$
$\mathbf{vp}_{t,e}=\mathbf{vp}_{i,c}=\mathbf{vp}_{i,s}=\mathbf{vp}_{e},$
$\mathbf{vp}_{t,c}=\mathbf{vp}_{c},\mathbf{vp}_{t,s}=\mathbf{vp}_{s},$
$\dot{\mathbf{vp}}_{t,c}=\ddot{\mathbf{vp}}_{t,c}=\dot{\mathbf{vp}}_{t,s}=\ddot{\mathbf{vp}}_{t,s}=\{\mathbf{0},0\}.$ (7)
The rest of the parameters are optimizable, and these parameters are shown in Fig. 3 (green). To ensure smoothness, three trajectories share the same velocity and acceleration at the connected point.
The objective function of the ASEO trajectory optimization problem is formulated as follows:
$\textrm{J}_{ASEO}(\mathcal{J}_{e},\mathcal{J}_{c},\mathcal{J}_{s})=j_{e,f}+j_{e,t}+j_{c,f}+j_{c,t}+j_{s,f},$ (8)
where $j_{e,f},j_{c,f}$, and $j_{s,f}$ are the feasibility costs of $\mathcal{J}_{e},\mathcal{J}_{c}$, and $\mathcal{J}_{s}$ corresponingly. To speed up the continuous exploration, $j_{e,t}$ and $j_{c,t}$ (the time costs of $\mathcal{J}_{e}$ and $\mathcal{J}_{c}$) are optimized. The feasibility cost $j_{\mu,f},\mu\in\{e,c,s\}$, it is computed by
$j_{\mu,f}=\int_{t_{s},\mu}^{t_{e,\mu}}\lambda_{p}\mathrm{Vp}_{\mu}(t))+\lambda_{d}(\mathrm{Vv}_{\mu}(t)+\mathrm{Va}_{\mu}(t))dt,$ (9)
where $\mathrm{Vp}_{\mu}(t)$ calculates the position violation of collision-free corridor constraint, $\mathrm{Vv}_{\mu}(t)$ calculates the linear velocity violation of maximum linear velocity constraint, and $\mathrm{Va}_{\mu}(t)$ calculates the maximum acceleration violation of linear acceleration constraint. The penalty weights are $\lambda_{p}$ and $\lambda_{d}$. For the time cost $j_{\mu,t},\mu\in\{e,c\}$, it is computed by
$j_{\mu,f}=\lambda_{t}\sum\limits_{i=1}^{m_{\mu}}\Delta t_{\mu,i},$ (10)
where $t_{\mu,i}\in\Delta T_{\mu}$, $m_{\mu}$ is the number of trajectory segments, and $\lambda_{t}$ is the penalty weight of time. For details of derivation and optimization, we refer the readers to [23].
After generating the ASEO trajectory, $\mathcal{J}_{e}$ is executed. Once $\mathbf{vp}_{e}$ is explored or obstacles are detected on $\mathcal{J}_{c}$, the total planner will plan again.
V experiments and results
V-A Implementation Details
To evaluate the proposed method, we compare our method against state-of-the-art methods [20, 16] in simulation experiments, and conduct real-world experiments.
For comparative simulations, all the methods are implemented in C++ on a computer that runs the Ubuntu 20.04 LTS operation system, it uses Robot Operation System (ROS) for communication between the planner and simulator. The computer is with an Intel Core i7-12700H, a GeForce 3060 12G, and 16G memory. For a fair comparison, all the methods are simulated on the platform that is provided by [20]. The FOV of the RGB-D camera used in the simulation is $[115,92]^{\circ}$, and its maximum sensing range is $5.0\,m$. As for real-world experiments, the proposed method is deployed on a self-made UAV platform (Fig.6 (a)). The proposed method is run on an on-board computer with an Intel Core i7-1260P. The UAV is equipped with a Livox Mid 360 LiDAR for localization and an Intel Realsense D435i RGB-D camera for exploration. The FOV of the Realsense camera is $[87,58]^{\circ}$, and its maximum sensing range is $5.0\,m$. The parameters of our method are shown in Table. I.
TABLE I: Parameters of the proposed method
$r_{v}=0.1m$ $v_{m}=2.0\,m/s$ $a_{m}=3.0\,m/s^{2}$ $\dot{\psi}_{m}=1.57\,rad/s$
$\ddot{\psi}_{m}=1.57\,rad/s^{2}$ $\lambda_{a}=0.6$ $\lambda_{arc}=1.0$ $\lambda_{e}=0.3$
$\lambda_{p}=4000$ $\lambda_{d}=350$ $\lambda_{t}=700$
V-B Simulation Results
In this section, we compare the proposed method against [20, 16] in four simulation environments. Each method runs 10 times in a simulation environment. We record the exploration time, UAV speed, traveled distance, and computational cost for evaluation (shown in Table. II).
  • FALCON [20]: This method plans the exploration sequence of unexplored regions by solving ATSP. After planning the exploration sequence, a time-optimal B-spline trajectory is generated utilizing [38].
  • FUEL [16]: This method plans the exploration sequence of frontiers through solving ATSP. It also uses time-optimal B-spline trajectories for exploration.
TABLE II: EXPLORATION STATISTIC OF ALL METHODS
Environment Method Exploration Time (s) Computational Cost ($ms$) UAV Speed ($m/s$) Traveled Distance ($m$)
Avg Std Avg Std Avg Std Avg Std
DARPA Tunnel Proposed 53.48 1.24 10.31 2.25 1.80 0.018 93.38 2.05
FALCON 56.24 4.50 19.80 6.13 1.63 0.039 96.88 5.84
FUEL 65.32 2.18 10.19 4.59 1.40 0.027 91.78 3.83
Classical Office Proposed 91.56 3.60 14.42 2.94 1.79 0.016 160.86 6.64
FALCON 102.68 3.52 20.00 10.11 1.59 0.046 155.25 6.45
FUEL 128.56 5.71 22.76 18.74 1.31 0.045 167.89 10.65
Large Maze Proposed 301.64 15.65 16.16 3.77 1.86 0.015 557.07 32.13
FALCON 343.84 12.45 63.38 33.03 1.61 0.018 554.98 21.64
FUEL 430.08 13.65 97.81 76.42 1.41 0.011 605.55 25.51
Large Tunnel Proposed 2441.32 51.78 20.48 5.70 3.32 0.017 8095.60 185.38
FALCON - - - - - - - -
FUEL - - - - - - - -
(a) Proposed.
(b) FALCON.
(c) FUEL.
(d) Explored Volume.
(e) Proposed.
(f) FALCON.
(g) FUEL.
(h) Explored Volume.
(i) Velocity Color.
(j) Proposed.
(k) FALCON.
(l) FUEL.
(m) Explored Volume.
Figure 4: The mapping results and the executed trajectory of all the methods. The color of the trajectory denotes the velocity of the UAV. Each row illustrates the simulations in one environment. The right column illustrates the explored volume vs. time.
The first simulation environment is the DARPA Tunnel with a scale of $42\times 20\times 2\,m^{3}$. Each method runs $75\,s$. This environment is relatively simple, so the main factor that affects the exploration efficiency is the speed of the UAV. As shown in Fig.4 (a-c), the proposed method maintains a higher speed than other methods ($10.4\%$ faster than FALCON and $28.6\%$ faster than FUEL). As a result, the proposed method is $5.2\%$ more efficient than FALCON and $16.1\%$ more efficient than FUEL.
The second simulation environment is the Classical Office with a scale of $15\times 30\times 2\,m^{3}$. Each method runs $130\,s$. Compared with DARPA Tunnel, Classical Office is more complex. The walls force the UAV to keep turning which slows down the speed of the UAV. Thanks to the curvature-penalized viewpoints planning and ASEO trajectory, the proposed method is able to avoid sharp turns and generate high-speed trajectories as shown in Fig. 4 (e-g). The UAV speed of the proposed method is $12.6\%$ faster than FALCON’s and $36.6\%$ faster than FUEL’s. Since this environment is more open than DARPA Tunnel, the computational costs grow. The computational cost proposed method is $28.9\%$ lower than FALCON’s and $36.6\%$ lower than FUEL’s. As a result, the proposed method is $12.1\%$ more efficient than FALCON and $40.4\%$ more efficient than FUEL.
The third simulation environment is the Maze with a scale of $40\times 40\times 2\,m^{3}$. Each method runs $500\,s$. As shown by Fig. 4 (j-l), the proposed method keeps high speed in this environment. Because this environment is larger than previous ones, the computational costs of FALCON and FUEL are much higher. The computational cost of the proposed method also grows, but only $12.1\%$ compared to simulation in Classical Office. This is because EOCH can solve ATSP efficiently. The path lengths of the proposed method and FALCON remain similar, which are shorter than FUEL’s, though FALCON and FUEL have better solutions for ATSP. This is because their implicit optimal assumptions fail most of the time. Thus, finding a perfect ATSP solution in exploration is usually not cost-effective. The results show the proposed method’s exploration efficiency is $14.0\%$ more efficient than FALCON and $42.6\%$ more efficient than FUEL.
The last simulation environment is the Large Tunnel with a scale of $251.5\times 327.5\times 2\,m^{3}$. Since this environment is very large, we adopt $v_{m}=4.0m/s$ and $a_{m}=6.0m/s^{2}$ to save the simulation time. The purpose of this trial is to test the computational cost of the proposed method in an extremely large environment. The comparative method FALCON and FUEL fail to run in this environment due to their memory usage exceeding the maximum 16G memory of the computer. On the contrary, the proposed method can accomplish the exploration as shown by Fig. 5. The average computational cost of planning is $20.48\,ms$ which is low enough to plan in real time. The average speed is $3.32\,m/s$. The average exploration time is $2441.32\,s$. The result shows the proposed method is able to explore large-scale environments.
(a) Executed Trajectory and Mapping Result.
(b) Exploration Volume.
Figure 5: Simulation result in Large Tunnel.
(a) UAV Platform.
(b) Executed Trajectory and Mapping Result.
(c) Outdoor Environment.
(d) Exploration Volume.
Figure 6: Experiment results of real-world exploration.
V-C Real-World Experiment
To evaluate the proposed method further, we deploy our method on a self-made UAV (Fig. 6(a)) to explore a $42\times 21\times 3\,m^{3}$ outdoor environment (Fig. 6(c)). The exploration results are shown in Fig. 6 (b) and (d). The exploration cost $145.1\,s$ to travel $260.9\,m$, explore $2520\,m^{3}$ of the environment is explored. During the exploration, the average velocity of the UAV is $1.79m/s$, and the average planning time is $18.5\,ms$.
The simulation results show that the proposed method can plan high-speed exploration trajectories in real-time even in large 3-D environments. The real-world experiments show that the proposed method is also effective.
VI conclusion
In this work, we propose an efficient dual-layer exploration planning method (EDEN). By finding long-term routings through the proposed EOHC algorithm, deciding short-term targets through curvature-penalized evaluation, and generating ASEO trajectories, the proposed method is able to explore large environments efficiently with low computational cost and high speed. The simulation indicates the exploration efficiency of the proposed method is $5.2\%-14.0\%$ higher than state-of-the-art methods. The computational cost of the proposed method is low enough to enable exploration in large-scale environments.

### v2

@@
 Abstract
-Efficient autonomous exploration in large-scale environments remains challenging due to the high planning computational cost and low-speed maneuvers.
-In this paper, we propose a fast and computation-efficient dual-layer exploration planning method.
-The insight of our dual-layer method is efficiently finding an acceptable long-term region routing and greedily exploring the target in the region of the first routing area with high speed.
-Specifically, the proposed method finds the long-term area routing through an approximate algorithm to ensure real-time planning in large-scale environments.
-Then, the viewpoint in the first routing region with the lowest curvature-penalized cost which can effectively reduce decelerations caused by sharp turn motions will be chosen as the next exploration target.
-To further speed up the exploration, we adopt an aggressive and safe exploration-oriented trajectory to enhance exploration continuity.
-The proposed method is compared to state-of-the-art methods in challenging simulation environments.
-The results show the proposed method outperforms other methods in terms of exploration efficiency, computational cost, and trajectory speed.
+Efficient autonomous exploration in large-scale environments remains challenging due to high planning computational cost and low-speed maneuvers.
+In this paper, we propose an Efficient Dual-layer Exploration plaNning method (EDEN).
+The insight of our dual-layer planning method is efficiently finding an acceptable long-term region routing and greedily exploring the target in the first region of routing at high speed.
+Specifically, we propose a long-term region routing approximate algorithm, called “exploration-oriented heuristic double-tree (EOHDT) algorithm”, to ensure real-time planning in large-scale environments.
+Then, the viewpoint in the first routing region with the highest curvature-penalized score, which can effectively reduce decelerations caused by sharp turn motions, will be chosen as the next exploration target.
+To further speed up the exploration, we propose an aggressive and safe exploration-oriented trajectory planning approach to enhance exploration continuity and speed.
+The proposed method is compared with state-of-the-art methods in challenging simulation environments.
+The results show that the proposed method outperforms other methods in terms of exploration efficiency, computational cost, and trajectory speed.
 We also conduct real-world experiments to validate the effectiveness of the proposed method.
 The code will be open-sourced [^1] https://github.com/NKU-MobFly-Robotics/EDEN.
-Index Terms:
-Autonomous exploration, motion planning, trajectory planning.
 I INTRODUCTION
-Umanned aerial vehicles (UAVs) are popular in many 3-D exploration relevant applications, e.g., search and rescue [1, 2], environment inspection [3, 4], environment reconstruction [5, 6, 7, 8], among others.
-However, most existing methods remain low efficiency, low trajectory speed, and high computational cost.
-Existing exploration methods can be coarsely separated into two kinds of methods: greedy methods and optimal-routing-based methods.
+Unmanned aerial vehicles (UAVs) are popular in many 3-D exploration-relevant applications, e.g., search and rescue [1, 2], environment inspection [3, 4, 5], environment reconstruction [6, 7, 8, 9], healthcare services [10, 11], among others.
+However, most existing methods suffer from low efficiency, low-speed trajectories, and high computational cost.
+Existing exploration methods can be coarsely separated into three kinds of methods: greedy methods, optimal-routing-based methods, and learning-based methods.
 Greedy methods attempt to select the target that maximizes the reward in a short period.
-The original frontier-based method [9] selects the closest frontier, i.e., the boundary voxel of explored and unknown space as an exploration target.
-To improve the UAV speed, [10] propose to select a frontier in camera field-of-view (FOV) to minimize the velocity change.
+The original frontier-based method [12] selects the closest frontier, i.e., the boundary voxel of explored and unknown space as the exploration target.
+To improve the UAV speed, the work in [13] proposes to select a frontier in the camera field-of-view (FOV) to minimize the velocity change.
 By combining next-best-views and rapid random tree (RRT), receding-horizon-next-best-views (RH-NBVP) [14] is proposed, and its variant methods [15, 16, 17, 18] show high exploration efficiency in small environments.
 However, these greedy methods suffer from massive back-and-forth maneuvers in large-scale environments due to the lack of long-term consideration.
-Another type of method, optimal-routing-based methods [16, 17, 18, 19, 20, 21], make decisions by constructing and solving routing problems to minimize the total routing length.
+The optimal-routing-based methods [19, 20, 21, 22, 23], make decisions by constructing and solving routing problems to minimize the total routing length.
 The work in [23] proposes to find a coverage path by solving the traveling salesman problem (TSP).
 The method proposed by [19] constructs and solves the asymmetrical traveling salesman problem (ATSP) to find the shortest routing of frontiers.
-To further reduce back-and-forth maneuvers, [18] increases the routing priority of isolated frontiers.
+To further reduce back-and-forth maneuvers, the approach in [20] increases the routing priority of isolated frontiers.
 By solving routing problems, these methods dramatically reduce back-and-forth maneuvers.
-However, the computational cost grows fast with the number of routing nodes, which makes these methods have to execute stops for planning in large-scale environments.
-What is more, the work in [22] shows that the speed of UAVs in exploration still has great potential to be stimulated.
-Therefore, a computation-efficient, far-sighted, and high-speed method is expected.
+However, the computational cost grows fast with the number of routing nodes, which makes these methods have to execute stops to wait for planning results in large-scale environments.
+In addition, the works in [24, 22] show that the speed of UAVs in exploration has great potential to be stimulated and can be significantly increased, while most existing state-of-the-art methods fail.
+Some learning-based methods [25, 26, 27] are proposed to explore the large and complex environments.
+However, they share the common limitation of suboptimal path planning due to discrete action spaces.
+Therefore, a computation-efficient, far-sighted, and high-speed exploration method is highly expected.
 Motivated by the above drawbacks of existing methods, we propose an Efficient Dual-layer Exploration plaNning method (EDEN) which can plan efficient exploration motions in hundred-meter-level 3-D environments in real-time.
-Different from greedy methods and optimal-routing-based methods, the proposed dual-layer method first finds a long-term sub-optimal exploration region routing efficiently and then decides the short-term targets that minimizes the exploration motion cost greedily.
-Specifically, we improve the Christofides algorithm, which is an approximate algorithm for solving TSP, to an exploration-oriented heuristic Christofides (EOHC) algorithm to find a sub-optimal but satisfying tour for ATSP with low computational cost.
-Then, a viewpoint for safety and two viewpoints with the highest curvature-penalized exploration gain in the first exploration region of the routing will be selected for high-speed trajectory planning.
-Then, inspired by [23], we propose an aggressive and safe exploration-oriented (ASEO) trajectory for high-speed continuous exploration.
+Different from greedy methods and optimal-routing-based methods, the proposed dual-layer method first finds a long-term sub-optimal region routing efficiently and then decides the short-term targets that can fully take advantage of the current UAV speed.
+Specifically, we propose the exploration-oriented heuristic double-tree (EOHDT) algorithm to find a sub-optimal satisfactory routing for ATSP with low computational cost.
+Then, a viewpoint for safety and two viewpoints with the highest curvature-penalized exploration score in the first region of the routing will be selected for high-speed trajectory planning.
+Then, inspired by [28], we propose an aggressive and safe exploration-oriented (ASEO) trajectory planning method for high-speed continuous exploration.
 The ASEO trajectory is composed of three parts: an exploring part, a continuous part, and a safety part.
-The time of the exploring part, the time of the continuous part, and the feasibility of the safety part are optimized together to find an aggresive trajectory and a backup trajectory for safety.
+The time of the exploring part, the time of the continuous part, and the feasibility of the safety part are optimized simultaneously to compute an aggressive trajectory for fast exploration and a backup trajectory for safety.
 We evaluate the proposed method by comparing it against state-of-the-art methods in various challenging environments.
-The results show that the proposed method has $5.2\%-14.0\%$ higher exploration efficiency, $10.4\%-15.5\%$ faster UAV speed, and $28.9\%-74.5\%$ lower computational cost.
+The results show that the proposed method has $5.2\%-89.7\%$ higher exploration efficiency, $10.4\%-69.7\%$ faster UAV speed, and $28.9\%-93.9\%$ lower computational cost.
 Further, the proposed method is deployed on a UAV and is tested in real-world environments to evaluate its effectiveness.
@@
   1. 1.
-    We propose an exploration-oriented heuristic Christofides (EOHC) algorithm for efficiently solving large-scale ATSP.
-The proposed EOHC can find a satisfactory region routing with a low computational cost.
-The real-time performance of EOHC can be guaranteed even in hundred-meter-level environments.
+    We propose an exploration-oriented heuristic double-tree (EOHDT) algorithm for efficiently solving large-scale ATSP.
+EOHDT computes a satisfactory region routing with real-time performance even in hundred-meter-level environments.
   2. 2.
-    To take advantage of the current UAV speed and enhance exploration continuity, we propose a heuristic curvature-penalized continuous gain for the short-term exploration targets choosing.
+    To achieve high-speed continuous exploration, we propose a curvature-penalized continuous score evaluation for the determination of short-term exploration targets.
   3. 3.
-    To make the UAV keep high-speed continuous exploration in unknown environments, we propose an ASEO trajectory planning method that ensures trajectory efficiency and safety at the same time.
+    We propose the ASEO trajectory planning method that ensures trajectory efficiency and safety simultaneously.
   4. 4.
-    Massive comparative simulations are conducted to evaluate the proposed method. We also test the proposed method by real-world experiments. We will release our source code.
+    Extensive comparative simulations are conducted to evaluate the proposed method. We also test the proposed method by real-world experiments. We will release our source code.
 II RELATED WORK
@@
 It selects the frontier voxel that is closest to the robot as the exploration target.
-To minimize the UAV velocity change, [10] selects the frontier voxel that is located at the boundary of camera FOV as the exploration target.
+To minimize the UAV velocity change, the work in [13] selects the frontier voxel that is located at the boundary of the camera FOV as the exploration target.
 Different from the greedy frontier-based methods that only consider the next step exploration, Bircher et al.
-[11] propose RH-NBVP which samples a set of exploration motion sequences and executes the most informative sequence’s first step and it shows high efficiency in short-term exploration.
+[14] proposes RH-NBVP, which samples a set of exploration motion sequences and executes the most informative sequence’s first step, and it shows high efficiency in short-term exploration.
 Nevertheless, in large-scale environments, it suffers from sample failure.
 To cope with this problem, Selin et al.
-[24] use frontier to guide sampling when a local sample fails.
-However, long-distance sampling is time-consuming and it does not guarantee success even with the guidance of frontier.
-To reduce the planning cost and improve long-distance exploration, roadgraphs, which can compress the complex explorable space information and tarversable space information into compact nodes and edges, are widely used in recent proposed greedy methods [12, 25, 14, 26, 27, 28, 29, 30].
-Consequently, roadgraphs enable greedy methods of efficient planning in large environments.
+[29] use frontiers to guide sampling when a local sample fails.
+However, long-distance sampling is time-consuming, and it does not guarantee success even with the guidance of the frontier.
+To reduce the planning cost and improve long-distance exploration, road graphs, which can compress the complex explorable space information and traversable space information into compact nodes and edges, are widely used in recent proposed greedy methods [15, 30, 17, 9, 31, 32, 33].
+Consequently, greedy methods are able to plan efficiently on road graphs in large environments.
 Although these methods show high efficiency in short-term exploration and low computational cost, they suffer from back-and-forth maneuvers due to a lack of long-term consideration.
@@
 Optimal-routing-based methods usually formulate the exploration target decision problems into routing problems.
-Optimal-tour-based methods can be roughly devided into frontier-ATSP methods [16, 17, 31, 18] and coverage-path-ATSP methods [32, 20, 21, 5, 19].
-Frontier-ATSP methods formulate ATSP using frontiers and find the shortest tour between frontiers.
-Coverage-path-ATSP methods formulate ATSP with unexplored space and find the shortest coverage path.
-The main drawback of the optimal-routing-based methods is that they have to spend luxury computation on solving ATSP in large-scale environments.
-Moreover, counterintuitively, recent works [28, 33] report that optimal-tour-based methods usually do not have an obvious advantage in minimizing the exploration path distance compared to greedy methods.
-This is because optimal-routing-based methods have implicit optimal assumptions, where frontier-ATSP methods assume a new frontier will not be generated after exploring an existing frontier, while coverage-path-ATSP methods assume all the unexplored space is traversable.
+Optimal-tour-based methods can be roughly devided into frontier-ATSP-based methods [19, 34, 20] and coverage-ATSP-based methods [35, 22, 23, 6, 21].
+Frontier-ATSP-based methods formulate ATSP using frontiers and find the shortest tour between frontiers.
+Coverage-ATSP-based methods formulate ATSP with unexplored space and find the shortest coverage path.
+The main drawback of the optimal-routing-based methods is that they have to spend an extremely high computational cost on solving ATSP in large-scale environments.
+Moreover, recent works [31, 36] report that optimal-routing-based methods usually do not have an obvious advantage in minimizing the exploration path distance compared to greedy methods.
+This is because optimal-routing-based methods usually make implicit optimal assumptions.
+Frontier-ATSP-based methods assume new frontiers will not arise after exploring a frontier.
+Coverage-ATSP-based methods assume all the unexplored space is traversable for the robot.
 Obviously, the above assumptions do not hold in most cases.
-Therefore, spending too much computation on finding a high-quality ATSP solution is not cost-effective.
-II-C Exploration-Oriented Trajectory Planning
-Exploration trajectory plays a crucial role in exploration tasks.
-The early proposed exploration methods [11, 9] stop-and-go motions are used.
-To utilize the good maneuverabilities of UAVs, Dharmadhikari et, al.
+Therefore, spending too much effort on computing a high-quality ATSP solution is not cost-effective.
+II-C Learning-Based Exploration Planning
+Learning-based methods have been used to tackle autonomous exploration problems with the advancement of machine learning.
+The work in [37] presents a modular deep-reinforcement-learning-based exploration framework that uses a convolutional Q-network to select goals from partial maps.
+The works in [38, 27] employ modular reinforcement learning frameworks for efficient visual exploration in unknown environments and provide systematic evaluations across realistic settings.
+The works in [26, 25] propose to construct a graph representation of the environment as the input to a graph neural network.
+However, they rely on discrete action spaces sampled from grid maps, which limits the optimality of the final path.
+In addition, the loading capacities of UAVs are usually limited, making them hard to carry extra devices for parallel computing.
+II-D Exploration-Oriented Trajectory Planning
+The exploration trajectory plays a crucial role in exploration tasks.
+The early proposed exploration methods [14, 12] use stop-and-go motions.
+To utilize the good maneuverabilities of UAVs, Dharmadhikari et al.
 [18] adopts motion primitives to speed up the exploration.
 To further improve the trajectory speed, Zhou et al.
-[16] optimizes minimum-time trajectory with optimizable terminal velocity and acceleration for high-speed exploration.
-Although minimum-time trajectory is very efficient, the optimizable terminal velocity and acceleration do not ensure being beneficial to continuously exploring the newly generated frontier after exploring the previous frontier.
-What is more, the optimizable terminal velocity and acceleration can lead to collision.
+[19] optimizes a minimum-time trajectory with optimizable terminal velocity and acceleration for high-speed exploration.
+Although the minimum-time trajectory is very efficient, its optimizable terminal velocity and acceleration do not ensure improving the exploration continuity when new frontiers are generated after exploring the old frontier.
+What is more, the optimizable terminal velocity and acceleration can lead to a collision without considering obstacles that may be detected in the unknown space.
 In summary, a far-sighted and computationally efficient exploration planning method and a trajectory planning method that enable fast continuous exploration while ensuring safety are required for exploration in large-scale environments.
@@
 III SYSTEM OVERVIEW
-The proposed method is composed of a dual-layer exploration target decision module that can plan both the long-term region routing and the short-term viewpoint target, and a trajectory planning module that can plan high-speed continuous exploration trajectories as shown in Fig.
-1.
-A dynamic voxel map with resolution $r_{v}$ is updated by integrating depth image and odometry information.
-Different from the normal voxel map, the dynamic voxel map only maintains a local part of voxels around the UAV while other voxels are dynamically read and written to save memory usage in large environments.
-Based on the voxel map, a dynamic topological graph (DTG), is maintained to manage exploration-related information.
-The traversable information is maintained by DTG’s history nodes and edges between them, and the explorable space information is maintained by DTG’s explorable region of interest (EROI), which is a set of uniform scale cubic regions that can be explored by visiting corresponding active viewpoints.
+The dual-layer exploration target decision module plans both the long-term region routing and the short-term viewpoint targets.
+The trajectory planning module is responsible for planning high-speed continuous exploration trajectories.
+To achieve fast decision-making, a dynamic voxel map with resolution $r_{v}$ is updated by integrating depth image and odometry information.
+Different from the normal voxel map, the dynamic voxel map only maintains a local part of voxels around the UAV, while other voxels are dynamically read and written to save memory usage in large environments.
+Based on the voxel map, a dynamic topological graph (DTG, Fig.
+2(a)) is maintained to manage exploration-related information.
+The traversable information is maintained by DTG’s history nodes and edges between them, and the explorable space information is maintained by DTG’s explorable region of interest (EROI), which is a set of uniform-scale cubic regions that can be explored by visiting corresponding active viewpoints.
 For detailed information, we refer readers to our previous work [39].
-The exploration planning utilizes DTG to efficiently formulate ATSP, and ATSP is solved through EOHC to find satisfactory exploration region routing with small computation.
-Then, three viewpoints, including an exploring viewpoint, a continuous viewpoint, and a safety viewpoint in the first region of the region routing are found for the following trajectory planning.
-Then, an aggressive and safe exploration-oriented trajectory which separates at the exploring viewpoint, and its two branches end at a continuous viewpoint and a safe viewpoint correspondingly, is optimized.
-The exploring part of the aggressive and safe exploration-oriented trajectory will be executed.
+The exploration planning utilizes DTG to efficiently formulate ATSP, and ATSP is solved through EOHDT to find satisfactory exploration region routing with small computation.
+Then, three viewpoints, i.e., an exploring viewpoint, a continuous viewpoint, and a safety viewpoint in the first region of the region routing, are selected for the following trajectory planning.
+With these three viewpoints, an aggressive and safe exploration-oriented trajectory is generated, which is divided into three parts from the exploring viewpoint.
+The first part is named the exploring part, which goes from the UAV’s current position to the exploring viewpoint.
+The second part is named the continuous part, which goes from the exploring viewpoint to the continuous viewpoint.
+The last part is named the safety part, which goes from the exploring viewpoint to the safety viewpoint.
+After every replanning, the exploring part of the aggressive and safe exploration-oriented trajectory will be actually executed.
 IV DUAL-LAYER EXPLORATION VIEWPOINTS PLANNING
-In this section, we first propose the EOHC algorithm for finding a satisfactory region routings efficiently in large-scale environments.
-Then, we use curvature-penalized continuous gain to decide the targets in the first region of the tour for the following high-speed ASEO trajectory planning.
-Figure 2: An instance of EOHC-based routing planning, the numbers indicate the lengths of edges.
-(a) An MST that connects all the explorable history nodes of DTG is extracted.
-(b) The leaf point, whose distance is the longest to the root node, and its corresponding branch (green) will be executed last.
-(c) Once, there are multiple siblings for the current node, the sibling (orange dashed curve) whose first node has the shortest distance to the current node will be chosen as the next branch.
-(d) The final routing found by the proposed EOHC-based routing planning.
-IV-A EOHC-Based Global Routing Planning
-As analyzed in Section.
-II.B, the optimal routing of frontiers or unexplored space is very likely not the optimal routing for exploration due to the implicit optimal assumption not being met.
-Therefore, instead of finding a high-quality of all precise explorable frontiers, the proposed method aims to spend low computation on finding a satisfactory coarse routing of explorable regions.
-The process of EOHC-based routing planning is shown in Fig.
-2.
-Firstly, we extract all the history nodes of DTG that connect at least one EROI as the explorable regions.
-To reduce the computation when formulating ATSP, the distances between all explorable regions and the distances between the robot and all explorable regions are computed by conducting the Dijkstra algorithm on DTG.
+In this section, we first propose the EOHDT algorithm for finding a satisfactory region routing efficiently in large-scale environments.
+Then, we use a curvature-penalized continuous score to decide the targets in the first region of the tour for the following high-speed ASEO trajectory planning.
+Figure 2: An illustration of DTG and an instance of EOHDT-based routing planning.
+(a) DTG and its boundary regions (pink dashed boxes).
+(b) An MST that connects all the boundary regions of DTG is extracted.
+(c) The leaf point, whose distance is the longest to the root node, and its corresponding branch (green) will be executed last.
+The numbers indicate the lengths of edges.
+(d) Once there are multiple child nodes for the current node, the child node (orange dashed curve) whose first node has the shortest distance to the current node will be chosen as the next branch.
+(e) The final routing found by the proposed EOHDT-based routing planning.
+IV-A EOHDT-Based Global Routing Planning
+As analyzed in Section II, a far-sighted and computationally efficient exploration planner for the large-scale exploration tasks is required.
+Therefore, instead of finding a high-quality routing of all precise explorable frontiers, the proposed method aims to spend low computation on finding a satisfactory coarse routing of boundary regions.
+We propose the EOHDT-based routing planning, which has a low computational complexity and a length guarantee in the worst case, for solving ATSP.
+The process of EOHDT-based routing planning is shown in Fig.
+2(b-e).
+Firstly, we extract all the boundary regions.
+Here, a boundary region is composed of a history node and a set of EROIs (at least one) connected to this history node as shown in Fig.
+2(a).
+To reduce the computation when formulating ATSP, the distances between all boundary regions and the distances between the robot and all boundary regions are computed by conducting the Dijkstra algorithm on DTG.
 Thanks to the sparse DTG, the ATSP can be formulated efficiently.
-To solve ATSP efficiently, we design an EOHC algorithm, which is improved from the Christofides algorithm [35] and tailored to ATSP.
-The key idea of the Christofides algorithm is to construct a minimum spanning tree (MST, Fig.
-2(a)) using the Prim algorithm and find a walk that goes around the MST.
-The solution routing of the Christofides algorithm ensures no longer than twice the optimal routing.
-However, the Christofides algorithm is designed for symmetric TSP.
+To solve ATSP efficiently, we design an EOHDT algorithm, which is improved from the double-tree algorithm [40] and tailored to ATSP.
+The key idea of the double-tree algorithm is to construct a minimum spanning tree (MST, Fig.
+2(b)) using the Prim algorithm and find a walk that goes around the MST.
+The solution routing of the double-tree algorithm ensures no longer than twice the optimal routing.
+However, the double-tree algorithm is designed for symmetric TSP.
 For exploration-oriented ATSP, the robot does not have to return to the original node as stated by [19].
 Therefore, to minimize the routing length, the leaf node that is the farthest from the robot and nodes on its corresponding branch (in Fig.
-2(b), the green nodes) will be executed last to save the highest cost of returning to the original node.
-Another problem is how to decide the sequence of other branches.
-Since all the regions will be explored finally, the proposed method chooses the next node closest to the current node heuristically when the current node has multiple siblings as shown in Fig.
-2(c).
+2(c), the green nodes) will be executed last to save the highest cost of returning to the original node.
+Another problem is how to decide the sequence of the other branches.
+Since all the regions will be explored finally, the proposed method chooses the next node closest to the current node heuristically when the current node has multiple child nodes, as shown in Fig.
+2(d).
 Finally, a routing for ATSP will be found as shown in Fig.
-2(d).
-For the worst case, the path length $l_{eohc}$ of the routing solved by the EOHC algorithm holds:
-$l_{eohc}\leq 2l^{*}+l_{m},$ (1)
-where $l^{*}$ is the optimal solution of the corresponding ATSP, and $l_{m}$ is the longest distance among all the distances between the robot and all the explorable history nodes.
-The computational complexity of the proposed EOHC is $\mathcal{O}(n^{3})$, where $n$ is the total number of nodes.
+2(e).
+Theorem 1.
+The EOHDT runs in $\mathcal{O}(n^{2})$ time, where $n$ denotes the number of the history nodes.
+Proof.
+EOHDT starts by generating an MST utilizing the Prim algorithm, which takes $\mathcal{O}(n^{2})$ time.
+Then, EOHDT finds the farthest leaf node among $n_{leaf}$ leaf nodes, which takes $\mathcal{O}(n_{leaf})$ time.
+Let $m$ be the number of nodes that have multiple child nodes in the MST.
+Then, EOHDT needs to conduct $m$ times sorting of child nodes.
+It results in $\mathcal{O}(\sum_{i=1}^{m}n_{i}\log n_{i})$ time complexity, where $n_{i}$ is the child nodes number of the $i$-th node among the $m$ nodes.
+Because $n_{i}<n$ and $2\leq n_{i}$, we have $\sum_{i=1}^{m}n_{i}\log n_{i}<n^{2}$.
+Therefore, EOHDT runs in $\mathcal{O}(n^{2}+n_{leaf}+\sum_{i=1}^{m}n_{i}\log n_{i})\approx\mathcal{O}(n^{2})$ time.
+Theorem 2.
+The length of the routing solved by EOHDT $l_{E}$ holds $l_{E}\leq 2l^{*}-l_{m}$, where $l^{*}$ is the length of the optimal solution of the ATSP, and $l_{m}$ is the longest distance among distances between the robot and all the nodes on the MST.
+Proof.
+Given the summation of all the edge lengths of the MST $l_{MST}$, it holds $l_{MST}\leq l^{*}$. It is easy to derive
+${\color[rgb]{0.0,0.0,0.0}l_{E}=2l_{MST}-l_{m}\leq 2l^{*}-l_{m}.}$ (1)
 IV-B Curvature-Penalized Viewpoints Planning
 After finding the long-term routing of regions, three viewpoints (an exploring viewpoint, a continuous viewpoint, and a safety viewpoint) in the first region of the routing will be selected for ASEO trajectory planning.
-The UAV will go to explore the exploring viewpoint (in the explored map) and maintain a speed and an acceleration conducive to explore the continuous viewpoint (in the explored map or unexplored map) if no obstacle is detected, otherwise it can return to the safety viewpoint (in the explored map).
+The UAV will go to explore the exploring viewpoint (in the explored map) and maintain a speed and an acceleration conducive to exploring the continuous viewpoint (in the explored map or unexplored map) if no obstacle is detected, otherwise it can return to the safety viewpoint (in the explored map).
 For detailed viewpoint evaluation, existing methods usually use the path distance and turning angle to avoid long traveling distances and sharp turns in a decoupled manner.
 However, the maximum velocity is limited by the curvature of the trajectory at a point according to the circular motion model.
-We compute the velocity $v_{s}(\mathbf{p})$ at $\mathbf{p}$ by
+The maximum velocity $v_{s}(\mathbf{p})$ at $\mathbf{p}$ is computed by
 $v_{s}(\mathbf{p})=\min\left(\sqrt{\frac{\lambda_{a}\cdot a_{m}}{K_{c}(\textbf{p})}},v_{m}\right),$ (2)
-where $v_{m}$ is the maximum linear velocity of the UAV, $a_{m}$ is the maximum linear acceleration of the UAV, $K_{c}(\mathbf{p})$ is the curvature of the trajectory at $\mathbf{p}$, and $\lambda_{a}\in(0.0,1.0)$ is an adjustable parameter.
-Based on this fact, we designed a curvature-penalized heuristic motion cost evaluation method to select the best exploring target and the best continuous target for high-speed exploration.
+where $v_{m}$ is the maximum linear velocity of the UAV, $a_{m}$ is the maximum linear acceleration of the UAV, $K_{c}(\mathbf{p})$ is the curvature of the trajectory at $\mathbf{p}$, and $\lambda_{a}\in(0.0,1.0)$ is a tunable parameter for reducing sharp turns with high accelerations.
+In the following viewpoints planning, a smaller $\lambda_{a}$ leads to fewer turns but may miss the opportunities to explore some corner areas.
+Based on this fact, we designed a curvature-penalized motion score evaluation method to select the best exploring target and the best continuous target for high-speed exploration.
 IV-B1 Curvature-Penalized Exploring Viewpoint and Continuous Viewpoint Planning
 An exploring viewpoint in the first region of routing and a continuous viewpoint located in the corresponding EROI of the exploring viewpoint will be selected for aggressive high-speed trajectory planning.
-First, a Dijkstra search starts from the current UAV position in the explored space is conducted to find the path to each candidate viewpoint in the first region of the routing.
-Then, for each path to a candidate viewpoint, an anchor point on the path which is $\lambda_{arc}\cdot v_{m}$ ($\lambda_{arc}$ is a tunable parameter) away from the UAV position is selected to formulate a virtual circular arc trajectory.
-The formulation of the $i$-th virtual circular arc trajectory to the $k$-th candidate viewpoint $\textbf{vp}_{k}=\left\{\mathbf{p}_{k},\psi_{k}\right\}\in\mathbb{R}^{3}\times SO(2)$ is constrained by the current UAV position $\mathbf{p}_{c}\in\mathbb{R}^{3}$, the current UAV speed $\dot{\mathbf{p}}_{c}\in\mathbb{R}^{3}$, and the anchor point $\mathbf{p}_{a,k}\in\mathbb{R}^{3}$.
-The center point of the circular arc $\mathbf{p}_{ac,k}\in\mathbb{R}^{3}$ and its curvature can be computed by
-$\mathbf{p}_{ac,k}=$
-[…]
-$K_{c,k}=\frac{1}{\|\mathbf{p}_{ac,k}-\mathbf{p}_{c}\|_{2}}.$ (3)
-The virtual velocity $v_{as,k}$ to $\textbf{vp}_{k}$ is penalized by the curvature $K_{c,k}$ and is computed by Eq. 2. The motion cost of going to this exploring target is computed by
-$\mathrm{CostE}(\mathbf{vp}_{k})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{c})}{v_{as,k}},\frac{\mathrm{YD}(\psi_{k},\psi_{c})}{\dot{\psi}_{m}}\right),$ (4)
-where the function $\mathrm{PL}(\mathbf{p}_{c}$, $\mathbf{p}_{k})$ computes the length of the path from $\mathbf{p}_{c}$ to $\mathbf{p}_{k}$, $\mathrm{YD}(\psi_{k},\psi_{c})$ the yaw difference between the current UAV yaw $\psi_{c}$ and $\psi_{k}$, and $\dot{\psi}_{m}$ is the maximum yaw velocity.
-Since the terminal velocity at $\mathbf{vp}_{k}$ is unknown, we continue to use the $v_{as,k}$ for computing the continuous motion cost from $\mathbf{vp}_{k}$ to the $j$-th viewpoint $\mathbf{vp}_{j}=\left\{\mathbf{p}_{j},\psi_{j}\right\}\in\mathbb{R}^{3}\times SO(2)$ inside the EROI of $\mathbf{vp}_{k}$
-$\mathrm{CostC}(\mathbf{vp}_{j})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{j})}{v_{as,k}},\frac{\mathrm{YD}(\psi_{j},\psi_{k})}{\dot{\psi}_{m}}\right).$ (5)
-Finnally, the gain of this viewpoint-pair is evaluated by
-$\mathrm{G}(\mathbf{vp}_{k},\mathbf{vp}_{j})=\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{k})}\cdot(1+\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{j})}).$ (6)
-Here, $\lambda_{e}$ is a tunable parameter.
-The larger $\lambda_{e}$ is, the more sensitive it is to the motion costs.
-The viewpoint-pair with the highest gain is selected as the exploring viewpoint $\mathbf{vp}_{e}$ and the continuous viewpoint $\mathbf{vp}_{c}$.
-It is worth noting that the voxels on the continuous line between $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ can not be occupied but free and unknown.
-This ensures the existence of the possibility for continuous exploration.
-To efficiently check the feasibilities of the viewpoint pairs, we propose a continuous line-checking method that follows the fast collision checking fashion [36] to avoid recalculation of indexes of voxels at overlapping positions of continuous lines.
-Specifically, our method generates a table of indexes of voxel that needs to be checked in advance and iterates the table when checking an EROI.
+First, a Dijkstra search, which starts from the current UAV position in the explored space, is conducted to find the path to each candidate viewpoint in the first region of the routing.
+Then, for each path to a candidate viewpoint, an anchor point on the path, which is $1.0\,s\times v_{m}$ length away from the UAV or the path terminal if the path is shorter than $1.0\,s\times v_{m}$, is selected to formulate a virtual circular arc trajectory.
+The formulation of a virtual circular arc trajectory to the $j$-th candidate viewpoint $\mathbf{vp}_{j}=\left\{\mathbf{p}_{j},\psi_{j}\right\}\in\mathbb{R}^{3}\times[-\pi,\pi)$ is constrained by the current UAV position $\mathbf{p}_{u}\in\mathbb{R}^{3}$, the current UAV speed $\dot{\mathbf{p}}_{u}\in\mathbb{R}^{3}$, and the anchor point $\mathbf{p}_{a,j}\in\mathbb{R}^{3}$.
+The curvature $K_{u,j}$ of the virtual circular arc trajectory can be computed by
+$\theta=\arccos$ $(\frac{\dot{\mathbf{p}}_{u}^{\mathrm{T}}\cdot(\mathbf{p}_{a,j}-\mathbf{p}_{u})}{\|\dot{\mathbf{p}}_{u}\|_{2}\cdot\|\mathbf{p}_{a,j}-\mathbf{p}_{u}\|_{2}}),$
+$K_{u,j}$ $={\color[rgb]{0.0,0.0,0.0}\frac{2\sin(\theta)}{\|\mathbf{p}_{a,j}-\mathbf{p}_{u}\|_{2}}.}$ (3)
+The virtual velocity $v_{as,j}$ to $\mathbf{vp}_{j}$ is penalized by the curvature $K_{u,j}$ and is computed by Eq. 2. The motion cost of going to this exploring target is computed by
+$\mathrm{CostE}(\mathbf{vp}_{j})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{j},\mathbf{p}_{u})}{v_{as,j}},\frac{|\mathrm{YD}(\psi_{j},\psi_{u})|}{\dot{\psi}_{m}}\right),$ (4)
+where the function $\mathrm{PL}(\mathbf{p}_{j}$, $\mathbf{p}_{u})$ computes the length of the path from $\mathbf{p}_{u}$ to $\mathbf{p}_{j}$, $\mathrm{YD}(\psi_{j},\psi_{u}):\left[-\pi,\pi\right)\times[-\pi,\pi)\mapsto[-\pi,\pi)$ computes the yaw difference between the current UAV yaw $\psi_{u}$ and $\psi_{j}$ by angle unwrapping, and $\dot{\psi}_{m}$ is the maximum yaw velocity.
+Since the terminal velocity at $\mathbf{vp}_{j}$ is unknown, we continue to use the $v_{as,j}$ for computing the continuous motion cost from $\mathbf{vp}_{j}$ to the $k$-th viewpoint $\mathbf{vp}_{k}=\left\{\mathbf{p}_{k},\psi_{k}\right\}\in\mathbb{R}^{3}\times[-\pi,\pi)$ inside the EROI of $\mathbf{vp}_{j}$
+$\mathrm{CostC}(\mathbf{vp}_{k})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{j})}{v_{as,j}},\frac{|\mathrm{YD}(\psi_{k},\psi_{j})|}{\dot{\psi}_{m}}\right).$ (5)
+Finally, the score for evaluating the viewpoint pair is computed by
+${\color[rgb]{0.0,0.0,0.0}\mathrm{G}(\mathbf{vp}_{k},\mathbf{vp}_{j})=\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{j})}\cdot(1+\mathrm{e}^{-\lambda_{e}\mathrm{CostC}(\mathbf{vp}_{k})}).}$ (6)
+Here, $\lambda_{e}$ is a positive tunable parameter.
+For a larger $\lambda_{e}$, the planner is inclined to choose the exploring viewpoint with a low $\mathrm{CostE}(\mathbf{vp}_{j})$.
+For a smaller $\lambda_{e}$, the planner is inclined to choose the viewpoint pair with a small total motion cost.
+The viewpoint-pair with the highest score is selected as the exploring viewpoint $\mathbf{vp}_{e}=\{\mathbf{p}_{e},\psi_{e}\}$ and the continuous viewpoint $\mathbf{vp}_{c}=\{\mathbf{p}_{c},\psi_{c}\}$.
+It is worth noting that the voxels on the line between $\mathbf{p}_{j}$ and $\mathbf{p}_{k}$ can not be occupied but are free and unknown.
+This ensures that $\mathbf{p}_{k}$ can be observed at $\mathbf{p}_{j}$, which makes it possible to explore continuously.
+To efficiently check the feasibility of the viewpoint pairs, we propose a continuous line-checking method that follows the fast collision checking fashion [41] to avoid recalculation of indices of voxels at overlapping positions of continuous lines.
+Specifically, our method generates a table of indices of voxels that need to be checked in advance and iterates through the table when checking an EROI.
 Once an iterative voxel is found occupied, its corresponding viewpoint pairs are confirmed infeasible and will not be considered.
@@
 After finding $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$, another viewpoint in the first region of routing will be selected.
-The safety viewpoint $\mathbf{vp}_{s}$ is the backup target if new obstacles are detected on the trajectory between $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$.
-To find $\mathbf{vp}_{c}$ efficiently, we directly conduct the Dijkstra algorithm starting from $\mathbf{vp}_{e}$, and the viewpoint that closest to $\mathbf{vp}_{e}$ will be selected as $\mathbf{vp}_{c}$.
+The safety viewpoint $\mathbf{vp}_{s}=\{\mathbf{p}_{s},\psi_{s}\}$ will be a feasible backup target, which ensures the UAV being able to return to the safe space, if new obstacles are detected on the trajectory between $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$.
+Without $\mathbf{vp}_{s}$, the UAV could be too fast to brake at $\mathbf{vp}_{e}$.
+Although we could find the optimal safety viewpoint through solving another ATSP starting from $\mathbf{vp}_{e}$, we directly select the viewpoint closest to $\mathbf{vp}_{e}$ as $\mathbf{vp}_{s}$ to save the computation.
+This safety viewpoint is found by conducting a Dijkstra search starting from $\mathbf{vp}_{e}$ inside the free space.
+Once a viewpoint is searched, $\mathbf{vp}_{s}$ is set as this viewpoint, and the Dijkstra search stops.
 Figure 3: An instance of ASEO trajectory.
 The initial position, velocity, and acceleration of $\mathcal{J}_{e}$ (violet) are fixed with the current UAV position, velocity, and acceleration (yellow).
-The continuous trajectory $\mathcal{J}_{c}$ (red) ends at $\mathbf{vp}_{c}$.
-The safety trajectory $\mathcal{J}_{s}$ (red) ends at $\mathbf{vp}_{s}$.
-The optimizable parameters, including intermediate 4-D points, trajectory durations, velocity, and acceleration at $\mathbf{vp}_{e}$, are illustrated in green.
-IV-C ASEO Trajectory Optimization
-Inspired by [23], we propose to optimize our 4-D aggressive and safe exploration-oriented (ASEO) trajectory which is based on MINCO trajectory [37].
-Different from the 3-D MINCO trajectory, the fourth dimension is the yaw angle, and we optimize it with the position trajectory together.
-A 4-D ASEO trajectory $\mathcal{J}(t):\left[t_{s},t_{e}\right]\mapsto\mathbb{R}^{3}\times SO(2)$ ($t_{s}$ and $t_{e}$ are the start time and end time of the trajectory) is paramerized by the initial state $\textbf{S}_{i}=\left\{\mathbf{vp}_{i},\dot{\mathbf{vp}}_{i},\ddot{\mathbf{vp}}_{i}\right\}$, the terminal state $\mathbf{S}_{t}=\left\{\mathbf{vp}_{t},\dot{\mathbf{vp}}_{t},\ddot{\mathbf{vp}}_{t}\right\}$, the $m-1$ intermediate viewpoints $\mathbf{VP}=\left\{\mathbf{vp}_{1},\mathbf{vp}_{2},\cdots,\mathbf{vp}_{m-1}\right\}$, and $m$ durations of trajectory segments $\Delta\mathbf{T}=\left\{\Delta t_{1},\Delta t_{2},\cdots,\Delta t_{m}\right\}$.
+The continuous trajectory $\mathcal{J}_{c}$ (red) ends at $\mathbf{vp}_{c}^{\prime}$.
+The safety trajectory $\mathcal{J}_{s}$ (blue) ends at $\mathbf{vp}_{s}^{\prime}$.
+The optimizable parameters, including intermediate 4-D points, trajectory durations, velocity, and acceleration at $\mathbf{vp}_{e}^{\prime}$, are illustrated in green.
+V ASEO TRAJECTORY OPTIMIZATION
+Inspired by the fast and safe trajectory planner (FASTER) [28], we propose to optimize our 4-D aggressive and safe exploration-oriented (ASEO) trajectory, which is based on the MINCO trajectory [42].
+Different from the 3-D MINCO trajectory, the fourth dimension is the yaw, and we optimize it with the position trajectory together.
+A 4-D MINCO trajectory $\mathcal{J}(t):\left[t_{s},t_{e}\right]\mapsto\mathbb{R}^{4}$ ($t_{s}$ and $t_{e}$ are the start time and end time of the trajectory) is paramerized by the initial state $\textbf{S}_{i}=\left\{\mathbf{vp}_{i},\dot{\mathbf{vp}}_{i},\ddot{\mathbf{vp}}_{i}\right\}$, the terminal state $\mathbf{S}_{t}=\left\{\mathbf{vp}_{t},\dot{\mathbf{vp}}_{t},\ddot{\mathbf{vp}}_{t}\right\}$, the m1m-1 intermediate viewpoints $\mathbf{VP}=\left\{\mathbf{vp}_{1},\mathbf{vp}_{2},\cdots,\mathbf{vp}_{m-1}\right\}$, and $m$ durations of trajectory segments $\Delta\mathbf{T}=\left\{\Delta t_{1},\Delta t_{2},\cdots,\Delta t_{m}\right\}$.
 As shown by Fig.
-3, the proposed ASEO trajectory is composed of three connected MINCO trajectories: an exploring trajectory $\mathcal{J}_{e}$ (paramerized by $\textbf{S}_{i,e}$, $\textbf{S}_{t,e}$, $\mathbf{VP}_{e}$, and $\Delta\mathbf{T}_{e}$), a continuous trajectory $\mathcal{J}_{c}$ (paramerized by $\textbf{S}_{i,c}$, $\textbf{S}_{t,c}$, $\mathbf{VP}_{c}$, and $\Delta\mathbf{T}_{c}$), and a safety trajectory $\mathcal{J}_{s}$ (paramerized by $\textbf{S}_{i,s}$, $\textbf{S}_{t,s}$, $\mathbf{VP}_{s}$, and $\Delta\mathbf{T}_{s}$).
+3, the proposed ASEO trajectory is composed of three connected 4-D MINCO trajectories: an exploring trajectory $\mathcal{J}_{e}$ (paramerized by $\mathbf{S}_{i,e}$, $\mathbf{S}_{t,e}$, $\mathbf{VP}_{e}$, and $\Delta\mathbf{T}_{e}$), a continuous trajectory $\mathcal{J}_{c}$ (paramerized by $\mathbf{S}_{i,c}$, $\mathbf{S}_{t,c}$, $\mathbf{VP}_{c}$, and $\Delta\mathbf{T}_{c}$), and a safety trajectory $\mathcal{J}_{s}$ (paramerized by $\mathbf{S}_{i,s}$, $\mathbf{S}_{t,s}$, $\mathbf{VP}_{s}$, and $\Delta\mathbf{T}_{s}$).
+The three trajectories of ASEO are connected at the exploring viewpoint $\mathbf{vp}_{e}$ to ensure the coverage of unknown space, which is different from FASTER [28], whose connecting point of its “Whole” trajectory and “Safety” trajectory can be any feasible point on the “Whole” trajectory.
 Some parameters of the ASEO trajectory are fixed.
-The exploring trajectory starts from the current UAV position $\{\mathbf{p}_{c},\psi_{c}\}$, velocity $\{\dot{\mathbf{p}}_{c},\dot{\psi}_{c}\}$, and acceleration $\{\ddot{\mathbf{p}}_{c},\ddot{\psi}_{c}\}$.
-The three trajectories are connected at the exploring viewpoint $\mathbf{vp}_{e}$.
-The continuous trajectory and the safety trajectory end at $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$ correspondingly, their velocity and acceleration are set by zero.
+The initial state of the exploring trajectory $\mathbf{S}_{i,e}=\left\{\mathbf{vp}_{i,e},\dot{\mathbf{vp}}_{i,e},\ddot{\mathbf{vp}}_{i,e}\right\}$ is aligned with the UAV position $\{\mathbf{p}_{c},\psi_{c}\}$, velocity $\{\dot{\mathbf{p}}_{c},\dot{\psi}_{c}\}$, and acceleration $\{\ddot{\mathbf{p}}_{c},\ddot{\psi}_{c}\}$.
+$\mathcal{J}_{e}$’s terminal pose $\mathbf{vp}_{t,e}$, $\mathcal{J}_{c}$’s initial pose $\mathbf{vp}_{i,c}$, and $\mathcal{J}_{s}$’s initial pose $\mathbf{vp}_{i,s}$ are fixed to the exploring viewpoint.
+$\mathcal{J}_{c}$’s terminal pose $\mathbf{vp}_{t,c}$ and $\mathcal{J}_{s}$’s terminal pose $\mathbf{vp}_{t,s}$ are fixed to the continuous viewpoint and the safety viewpoint correspondingly, their terminal velocity and acceleration are set to zero.
 We summarize these fixed parameters as:
 $\mathbf{vp}_{i,e},\dot{\mathbf{vp}}_{i,e},\ddot{\mathbf{vp}}_{i,e}=\{\mathbf{p}_{c},\psi_{c}\},\{\dot{\mathbf{p}}_{c},\dot{\psi}_{c}\},\{\ddot{\mathbf{p}}_{c},\ddot{\psi}_{c}\},$
-$\mathbf{vp}_{t,e}=\mathbf{vp}_{i,c}=\mathbf{vp}_{i,s}=\mathbf{vp}_{e},$
-$\mathbf{vp}_{t,c}=\mathbf{vp}_{c},\mathbf{vp}_{t,s}=\mathbf{vp}_{s},$
+$\mathbf{vp}_{t,e}=\mathbf{vp}_{i,c}=\mathbf{vp}_{i,s}=\mathbf{vp}_{e}^{\prime}=\{\mathbf{p}_{e},\psi_{e}^{\prime}\},$
+$\mathbf{vp}_{t,c}=\mathbf{vp}_{c}^{\prime}=\{\mathbf{p}_{c},\psi_{c}^{\prime}\},\mathbf{vp}_{t,s}=\mathbf{vp}_{s}^{\prime}=\{\mathbf{p}_{s},\psi_{s}^{\prime}\},$
 $\dot{\mathbf{vp}}_{t,c}=\ddot{\mathbf{vp}}_{t,c}=\dot{\mathbf{vp}}_{t,s}=\ddot{\mathbf{vp}}_{t,s}=\{\mathbf{0},0\}.$ (7)
-The rest of the parameters are optimizable, and these parameters are shown in Fig. 3 (green). To ensure smoothness, three trajectories share the same velocity and acceleration at the connected point.
+The rest of the parameters are optimizable, and these parameters are shown in Fig.
+3 (green).
+To ensure smoothness, three trajectories share the same velocity and acceleration at the connected point.
+It is worth noting that we use $\psi_{e}^{\prime}$, $\psi_{c}^{\prime}$, $\psi_{s}^{\prime}\in\mathbb{R}$ instead of $\psi_{e}$, $\psi_{c}$, $\psi_{s}\in[-\pi,\pi)$ to avoid the rotations greater than $\pi$.
+Specifically, $\psi_{e}^{\prime}$, $\psi_{c}^{\prime}$, and $\psi_{s}^{\prime}$ are calculated by $\psi_{e}^{\prime}=\mathrm{YD}(\psi_{e},\psi_{u})+\psi_{u}$, $\psi_{c}^{\prime}=\mathrm{YD}(\psi_{c},\psi_{e})+\psi_{e}^{\prime}$, and $\psi_{s}^{\prime}=\mathrm{YD}(\psi_{s},\psi_{e})+\psi_{e}^{\prime}$.
+The intermediate yaws of $\mathcal{J}_{e}$ are bounded within $\left[\min(\psi_{e}^{\prime},\psi_{u})-\pi,\max(\psi_{e}^{\prime},\psi_{u})+\pi\right]$, the intermediate yaws of $\mathcal{J}_{c}$ are bounded within $\left[\min(\psi_{c}^{\prime},\psi_{e}^{\prime})-\pi,\max(\psi_{c}^{\prime},\psi_{e}^{\prime})+\pi\right]$, and the intermediate yaws of $\mathcal{J}_{s}$ are bounded within $\left[\min(\psi_{s}^{\prime},\psi_{e}^{\prime})-\pi,\max(\psi_{s}^{\prime},\psi_{e}^{\prime})+\pi\right]$.
+Unlike FASTER, which optimizes different trajectories in a decoupled manner, we optimize the three trajectories of the ASEO trajectory simultaneously to ensure that each trajectory is feasible.
 The objective function of the ASEO trajectory optimization problem is formulated as follows:
 $\textrm{J}_{ASEO}(\mathcal{J}_{e},\mathcal{J}_{c},\mathcal{J}_{s})=j_{e,f}+j_{e,t}+j_{c,f}+j_{c,t}+j_{s,f},$ (8)
-where $j_{e,f},j_{c,f}$, and $j_{s,f}$ are the feasibility costs of $\mathcal{J}_{e},\mathcal{J}_{c}$, and $\mathcal{J}_{s}$ corresponingly.
+where $j_{e,f},j_{c,f}$, and $j_{s,f}$ are the feasibility costs of $\mathcal{J}_{e},\mathcal{J}_{c}$, and $\mathcal{J}_{s}$ correspondingly.
 To speed up the continuous exploration, $j_{e,t}$ and $j_{c,t}$ (the time costs of $\mathcal{J}_{e}$ and $\mathcal{J}_{c}$) are optimized.
-The feasibility cost $j_{\mu,f},\mu\in\{e,c,s\}$, it is computed by
-$j_{\mu,f}=\int_{t_{s},\mu}^{t_{e,\mu}}\lambda_{p}\mathrm{Vp}_{\mu}(t))+\lambda_{d}(\mathrm{Vv}_{\mu}(t)+\mathrm{Va}_{\mu}(t))dt,$ (9)
-where $\mathrm{Vp}_{\mu}(t)$ calculates the position violation of collision-free corridor constraint, $\mathrm{Vv}_{\mu}(t)$ calculates the linear velocity violation of maximum linear velocity constraint, and $\mathrm{Va}_{\mu}(t)$ calculates the maximum acceleration violation of linear acceleration constraint.
+The feasibility cost $j_{\mu,f},\mu\in\{e,c,s\}$ is computed by
+$j_{\mu,f}$ $=\int_{t_{s},\mu}^{t_{e,\mu}}\lambda_{p}\mathrm{Vp}_{\mu}(t)+\lambda_{d}(\mathrm{Vv}_{\mu}(t)+\mathrm{Va}_{\mu}(t)$
+$+\mathrm{V\dot{\psi}}_{\mu}(t)+\mathrm{V\ddot{\psi}}_{\mu}(t))dt,$ (9)
+where $\mathrm{Vp}_{\mu}(t)$ calculates the position violation of collision-free corridor constraint, and $\mathrm{Vv}_{\mu}(t)$, $\mathrm{Va}_{\mu}(t)$, $\mathrm{V\dot{\psi}}_{\mu}(t)$, and $\mathrm{V\ddot{\psi}}_{\mu}(t)$ calculate the violations of linear velocity, linear acceleration, yaw velocity, and yaw acceleration, respectively.
 The penalty weights are $\lambda_{p}$ and $\lambda_{d}$.
 For the time cost $j_{\mu,t},\mu\in\{e,c\}$, it is computed by
-$j_{\mu,f}=\lambda_{t}\sum\limits_{i=1}^{m_{\mu}}\Delta t_{\mu,i},$ (10)
-where $t_{\mu,i}\in\Delta T_{\mu}$, $m_{\mu}$ is the number of trajectory segments, and $\lambda_{t}$ is the penalty weight of time.
+${\color[rgb]{0.0,0.0,0.0}j_{\mu,t}=\lambda_{t}\sum\limits_{i=1}^{m_{\mu}}\Delta t_{\mu,i},}$ (10)
+where $\Delta t_{\mu,i}\in\Delta\mathbf{T}_{\mu}$, $m_{\mu}$ is the number of trajectory segments, and $\lambda_{t}$ is the penalty weight of time.
 For details of derivation and optimization, we refer the readers to [42].
-After generating the ASEO trajectory, $\mathcal{J}_{e}$ is executed. Once $\mathbf{vp}_{e}$ is explored or obstacles are detected on $\mathcal{J}_{c}$, the total planner will plan again.
-V experiments and results
-V-A Implementation Details
-To evaluate the proposed method, we compare our method against state-of-the-art methods [20, 16] in simulation experiments, and conduct real-world experiments.
-For comparative simulations, all the methods are implemented in C++ on a computer that runs the Ubuntu 20.04 LTS operation system, it uses Robot Operation System (ROS) for communication between the planner and simulator.
-The computer is with an Intel Core i7-12700H, a GeForce 3060 12G, and 16G memory.
+After generating the ASEO trajectory, $\mathcal{J}_{e}$ is executed.
+Our exploration planner replans (replans viewpoints and regenerates ASEO trajectory) when it encounters the following conditions: (1) the distance from an obstacle to a point on $\mathcal{J}_{e}$, $\mathcal{J}_{c}$, or $\mathcal{J}_{s}$ is shorter than the radius of the UAV $r_{u}$; (2) $\mathbf{vp}_{e}$ does not cover any unknown space or the UAV is about to reach $\mathbf{vp}_{e}$ ($\|\mathbf{p}_{u}-\mathbf{p}_{e}\|_{2}<d_{th}$ and $|\mathrm{YD}(\psi_{e},\psi_{u})|<\psi_{th}$, where $d_{th}$ is the distance threshold and $\psi_{th}$ is the yaw threshold); (3) the UAV has executed $\mathcal{J}_{e}$ for a replanning period $t_{r}$ or $\mathcal{J}_{e}$ will expire within a replanning offset time $t_{o}$.
+VI EXPERIMENTS and RESULTS
+VI-A Implementation Details
+To evaluate the proposed method, we compare our method against state-of-the-art methods [22, 19] in simulation experiments and conduct real-world experiments.
+For comparative simulations, all the methods are implemented in C++ on a computer that runs the Ubuntu 20.04 LTS operation system, which uses Robot Operation System (ROS) for communication between the planner and simulator.
+The computer has an Intel Core i7-12700H, a GeForce 3060 12G, and 16G memory.
 For a fair comparison, all the methods are simulated on the platform that is provided by [22].
-The FOV of the RGB-D camera used in the simulation is $[115,92]^{\circ}$, and its maximum sensing range is $5.0\,m$.
+The FOV of the RGB-D camera used in the simulation is $115,92]^{\circ}$, and its maximum sensing range is 5.0m5.0\,m.
 As for real-world experiments, the proposed method is deployed on a self-made UAV platform (Fig.6 (a)).
 The proposed method is run on an on-board computer with an Intel Core i7-1260P.
-The UAV is equipped with a Livox Mid 360 LiDAR for localization and an Intel Realsense D435i RGB-D camera for exploration.
-The FOV of the Realsense camera is $[87,58]^{\circ}$, and its maximum sensing range is $5.0\,m$.
+The UAV is equipped with a Livox Mid 360 LiDAR for localization (using Point-LIO [43]) and an Intel Realsense D435i RGB-D camera for exploration.
+The FOV of the Realsense camera is $87,58]^{\circ}$, and its maximum sensing range is 5.0m5.0\,m.
 The parameters of our method are shown in Table.
@@
 $r_{v}=0.1m$ $v_{m}=2.0\,m/s$ $a_{m}=3.0\,m/s^{2}$ $\dot{\psi}_{m}=1.57\,rad/s$
-$\ddot{\psi}_{m}=1.57\,rad/s^{2}$ $\lambda_{a}=0.6$ $\lambda_{arc}=1.0$ $\lambda_{e}=0.3$
-$\lambda_{p}=4000$ $\lambda_{d}=350$ $\lambda_{t}=700$
-V-B Simulation Results
-In this section, we compare the proposed method against [20, 16] in four simulation environments.
-Each method runs 10 times in a simulation environment.
-We record the exploration time, UAV speed, traveled distance, and computational cost for evaluation (shown in Table.
-II).
-  • FALCON [20]: This method plans the exploration sequence of unexplored regions by solving ATSP.
-After planning the exploration sequence, a time-optimal B-spline trajectory is generated utilizing [38].
-  • FUEL [16]: This method plans the exploration sequence of frontiers through solving ATSP. It also uses time-optimal B-spline trajectories for exploration.
+$\ddot{\psi}_{m}=1.57\,rad/s^{2}$ $\lambda_{a}=0.6$ $\lambda_{e}=0.3$ $\lambda_{p}=4000$
+$\lambda_{d}=350$ $\lambda_{t}=700$ $r_{u}=0.35\,m$ $d_{th}=0.6\,m$
+$\psi_{th}=0.25\,rad$ $t_{r}=2.0\,s$ $t_{o}=0.1\,s$
+(a) Proposed.
+(b) FALCON.
+(c) FUEL.
+(d) Proposed.
+(e) FALCON.
+(f) FUEL.
+(g) Proposed.
+(h) FALCON.
+(i) FUEL.
+(j) Velocity.
+(k) Proposed.
+(l) FALCON.
+(m) FUEL.
+(n) Explored Volume (DARPA Tunnel).
+(o) Explored Volume (Classical Office).
+(p) Explored Volume (Large Maze).
+(q) Explored Volume (City).
+Figure 4: The mapping results and the executed trajectory of all the methods.
+The color of the trajectory (j) denotes the velocity of the UAV.
+Each row illustrates the simulations in one environment.
+The last row illustrates the explored volume vs.
+time.
 TABLE II: EXPLORATION STATISTIC OF ALL METHODS
@@
 FUEL 430.08 13.65 97.81 76.42 1.41 0.011 605.55 25.51
+City Proposed 515.88 20.31 36.29 11.98 1.85 0.018 980.67 38.77
+FALCON 978.48 52.95 752.49 498.05 1.09 0.056 1083.57 35.53
+FUEL 998.32 165.41 596.04 608.50 1.09 0.173 1081.11 100.53
 Large Tunnel Proposed 2441.32 51.78 20.48 5.70 3.32 0.017 8095.60 185.38
@@
 FUEL - - - - - - - -
-(a) Proposed.
-(b) FALCON.
-(c) FUEL.
-(d) Explored Volume.
-(e) Proposed.
-(f) FALCON.
-(g) FUEL.
-(h) Explored Volume.
-(i) Velocity Color.
-(j) Proposed.
-(k) FALCON.
-(l) FUEL.
-(m) Explored Volume.
-Figure 4: The mapping results and the executed trajectory of all the methods.
-The color of the trajectory denotes the velocity of the UAV.
-Each row illustrates the simulations in one environment.
-The right column illustrates the explored volume vs.
-time.
+VI-B Simulation Results
+In this section, we compare the proposed method against [22, 19] in four simulation environments.
+For fair comparison, we modify the velocity and acceleration optimizations of the comparative methods.
+The modification constrains their trajectory generation within the maximum linear velocity and acceleration, rather than within the three axes individually.
+Except for the size of the exploration space, the other parameters of the comparative methods are not modified.
+Each method runs 10 times in a simulation environment.
+We record the exploration time, UAV speed, traveled distance, and computational cost for evaluation (shown in Table.
+II).
+  • FALCON [22]: This method plans the exploration sequence of unexplored regions by solving coverage-ATSP.
+After planning the exploration sequence, a time-optimal B-spline trajectory is generated utilizing [44].
+  • FUEL [19]: This method plans the exploration sequence of frontiers through solving frontier-ATSP. It also uses time-optimal B-spline trajectories for exploration.
 The first simulation environment is the DARPA Tunnel with a scale of $42\times 20\times 2\,m^{3}$.
-Each method runs $75\,s$.
+Each method runs 75s75\,s.
 This environment is relatively simple, so the main factor that affects the exploration efficiency is the speed of the UAV.
@@
 As a result, the proposed method is $5.2\%$ more efficient than FALCON and $16.1\%$ more efficient than FUEL.
+Fig.
+4(n) compares the explored volume of the three methods.
 The second simulation environment is the Classical Office with a scale of $15\times 30\times 2\,m^{3}$.
-Each method runs $130\,s$.
+Each method runs 130s130\,s.
 Compared with DARPA Tunnel, Classical Office is more complex.
-The walls force the UAV to keep turning which slows down the speed of the UAV.
-Thanks to the curvature-penalized viewpoints planning and ASEO trajectory, the proposed method is able to avoid sharp turns and generate high-speed trajectories as shown in Fig.
-4 (e-g).
+The walls force the UAV to keep turning, which slows down the speed of the UAV.
+Thanks to the curvature-penalized viewpoints planning and ASEO trajectory, the proposed method is able to reduce sharp turns and generate high-speed trajectories, while FALCON usually decelerates when meeting walls, and FUEL can reach high speed only in explored space, as shown in Fig.
+4(d-f).
+This is because the B-spline trajectories of FALCON and FUEL only try to minimize the current trajectory duration without considering continuous exploration.
+Moreover, the exploration efficiencies of FALCON and FUEL are limited because their implicit optimal assumptions are usually not fulfilled due to the wall obstacles and open space of this environment.
 The UAV speed of the proposed method is $12.6\%$ faster than FALCON’s and $36.6\%$ faster than FUEL’s.
-Since this environment is more open than DARPA Tunnel, the computational costs grow.
-The computational cost proposed method is $28.9\%$ lower than FALCON’s and $36.6\%$ lower than FUEL’s.
+Since this environment is more open than DARPA Tunnel, the computational costs of the three methods grow.
+The computational cost of the proposed method is $29.3\%$ lower than FALCON’s and $37.7\%$ lower than FUEL’s.
+The traveled distances of the three methods are similar.
 As a result, the proposed method is $12.1\%$ more efficient than FALCON and $40.4\%$ more efficient than FUEL.
+Fig.
+4(o) compares the explored volume of the three methods.
 The third simulation environment is the Maze with a scale of $40\times 40\times 2\,m^{3}$.
-Each method runs $500\,s$.
+Each method runs 500s500\,s.
 As shown by Fig.
-4 (j-l), the proposed method keeps high speed in this environment.
-Because this environment is larger than previous ones, the computational costs of FALCON and FUEL are much higher.
+4(g-i), the proposed method keeps a high speed in this environment.
+Because this environment is larger than previous ones, FALCON and FUEL have to spend more computing time on solving larger-scale ATSP.
 The computational cost of the proposed method also grows, but only $12.1\%$ compared to simulation in Classical Office.
-This is because EOCH can solve ATSP efficiently.
-The path lengths of the proposed method and FALCON remain similar, which are shorter than FUEL’s, though FALCON and FUEL have better solutions for ATSP.
-This is because their implicit optimal assumptions fail most of the time.
-Thus, finding a perfect ATSP solution in exploration is usually not cost-effective.
-The results show the proposed method’s exploration efficiency is $14.0\%$ more efficient than FALCON and $42.6\%$ more efficient than FUEL.
+This is because EOHDT can solve ATSP efficiently.
+Similar to the result in Classical Office, the traveled distance of the proposed method is slightly longer than FALCON’s but shorter than FUEL’s.
+This is caused by two factors.
+First, the curvature-penalized viewpoint planning makes the proposed method tend to choose to explore the target that can keep high UAV speed, which may leave some isolated unexplored corners and lead to the back-and-forth movements.
+Second, the routing solved by the proposed EOHDT is sub-optimal.
+Even so, the sacrifice of the traveled distance trade offers a lower computational cost and a faster UAV speed.
+The fourth simulation environment is the City with a scale of $40\times 40\times 8.5\,m^{3}$.
+Each method runs 1200s1200\,s.
+Compared with previous environments, City has a larger size and contains denser obstacles, which pose significant challenges to the computational efficiency of each method.
+As shown by Fig.
+4(k-m), the proposed method executes a straighter trajectory and keeps a higher speed, while comparative methods’ trajectories are more curved and their speeds are lower.
+The average computation time for both FALCON and FUEL exceeds 0.5s0.5\,s, which has a significant impact on the exploration of real-time planning.
+In contrast, thanks to EOHDT, the proposed method is able to plan in real-time with massive boundary regions.
+Consequently, the proposed method is $89.7\%$ more efficient than FALCON and $93.5\%$ more efficient than FUEL.
+Moreover, the traveled distance of the proposed method is shorter than that of FALCON.
+This is because City contains dense building obstacles, which are far from FALCON’s implicit optimal assumption.
 The last simulation environment is the Large Tunnel with a scale of $251.5\times 327.5\times 2\,m^{3}$.
+Each method runs 3000s3000\,s.
 Since this environment is very large, we adopt $v_{m}=4.0m/s$ and $a_{m}=6.0m/s^{2}$ to save the simulation time.
 The purpose of this trial is to test the computational cost of the proposed method in an extremely large environment.
-The comparative method FALCON and FUEL fail to run in this environment due to their memory usage exceeding the maximum 16G memory of the computer.
+The comparative methods, FALCON and FUEL, fail to run in this environment due to their high memory usage and large computation.
 On the contrary, the proposed method can accomplish the exploration as shown by Fig.
 5.
-The average computational cost of planning is $20.48\,ms$ which is low enough to plan in real time.
-The average speed is $3.32\,m/s$.
-The average exploration time is $2441.32\,s$.
+The average computational cost of planning is 20.48ms20.48\,ms, which is low enough to plan in real time.
+The average speed of the UAV is 3.32m/s3.32\,m/s.
+The explored volume of the proposed method is shown in Fig.
+5(b).
+The average exploration time is 2441.32s2441.32\,s.
 The result shows the proposed method is able to explore large-scale environments.
 (a) Executed Trajectory and Mapping Result.
-(b) Exploration Volume.
+(b) Explored Volume.
 Figure 5: Simulation result in Large Tunnel.
 (a) UAV Platform.
-(b) Executed Trajectory and Mapping Result.
-(c) Outdoor Environment.
-(d) Exploration Volume.
-Figure 6: Experiment results of real-world exploration.
-V-C Real-World Experiment
+(b) Courtyard.
+(c) Park.
+(d) Mapping Result (Courtyard).
+(e) Mapping Result (Park).
+(f) Explored Volume (Courtyard).
+(g) Explored Volume (Park).
+Figure 6: Experiment results of real-world explorations.
+VI-C Real-World Experiment
 To evaluate the proposed method further, we deploy our method on a self-made UAV (Fig.
-6(a)) to explore a $42\times 21\times 3\,m^{3}$ outdoor environment (Fig.
-6(c)).
+6(a)), and test it in two outdoor environments.
+The first environment is the $Courtyard$ with a scale of $42\times 21\times 3\,m^{3}$ as shown by Fig.
+6(b).
+This environment is composed of trees and bushes.
 The exploration results are shown in Fig.
-6 (b) and (d).
-The exploration cost $145.1\,s$ to travel $260.9\,m$, explore $2520\,m^{3}$ of the environment is explored.
-During the exploration, the average velocity of the UAV is $1.79m/s$, and the average planning time is $18.5\,ms$.
-The simulation results show that the proposed method can plan high-speed exploration trajectories in real-time even in large 3-D environments.
+6(d) and Fig.
+6(f).
+The exploration takes 145.1s145.1\,s to travel 260.9m260.9\,m, explore $2520\,m^{3}$ of the environment.
+During the exploration, the average velocity of the UAV is 1.79m/s1.79\,m/s, and the average planning time is 18.5ms18.5\,ms.
+The second environment is the $Park$ with a scale of $30\times 25\times 3\,m^{3}$ as shown by Fig.
+6(c).
+This environment is composed of a pergola, uneven slopes, and cluttered trees, which is more challenging than the first environment.
+The exploration results are shown in Fig.
+6(e) and Fig.
+6(g).
+The exploration takes 145.5s145.5\,s to travel 256.4m256.4\,m, explore $2131\,m^{3}$ of the environment.
+The average velocity of the UAV is 1.77m/s1.77\,m/s, and the average planning time is 14.1ms14.1\,ms.
+Since this environment is more cluttered and smaller, the number of viewpoints to be evaluated and the number of nodes in ATSP are fewer, which reduces the UAV speed and the planning time.
+The results show the proposed method is able to plan a high-speed trajectory with limited computation resources online.
+The simulation results show that the proposed method can plan high-speed exploration trajectories in real-time, even in large 3-D environments.
 The real-world experiments show that the proposed method is also effective.
-VI conclusion
+VII CONCLUSION
 In this work, we propose an efficient dual-layer exploration planning method (EDEN).
-By finding long-term routings through the proposed EOHC algorithm, deciding short-term targets through curvature-penalized evaluation, and generating ASEO trajectories, the proposed method is able to explore large environments efficiently with low computational cost and high speed.
-The simulation indicates the exploration efficiency of the proposed method is $5.2\%-14.0\%$ higher than state-of-the-art methods.
-The computational cost of the proposed method is low enough to enable exploration in large-scale environments.
+By integrating the EOHDT algorithm for long-term region routing, the curvature-penalized viewpoint selection for short-term target planning, and the ASEO trajectory generation for high-speed continuous flight, EDEN significantly improves exploration performance.
+Experimental results show that EDEN achieves up to $89.7\%$ higher exploration efficiency, $10.4\%-69.7\%$ faster UAV speed, and over $28.9\%$ reduction in computational cost compared to state-of-the-art methods.
+These advantages were consistently observed across multiple complex simulation environments and validated in real-world outdoor tests.
+The method enables real-time planning in large-scale environments while maintaining high-speed flight and safety.
+One drawback of the proposed method is that continuity-enhanced exploration planning increases the traveled distance.
+We plan to prioritize the exploration of isolated unknown regions to reduce redundant movements.